### v1

# font=small

Abstract
In this paper, we study the effects of fill probabilities and adverse fills on the trading strategy simulation process. We specifically focus on a stochastic optimal control market-making problem and test the strategy on ES (E-mini S&P 500), NQ (E-mini Nasdaq 100), CL (Crude Oil) and ZN (10-Year Treasury Note), which are some of the most liquid futures contract listed on the CME (Chicago Mercantile Exchange). We provide empirical evidence which shows how fill probabilities and adverse fills can significantly effect performance, and propose a more prudent simulation framework for dealing with this. Many previous works aim to measure different types of adverse selection in the limit order book, however, they often simulate price processes and market orders independently. This has the ability to largely inflate the performance of a short-term style trading strategy. Our studies show that using more realistic fill probabilities, and tracking adverse fills, in the strategy simulation process, more accurately portrays how these types of trading strategies would perform in reality.
Keywords: Algorithmic and High-Frequency Trading, Market-Making, Adverse Selection, Stochastic Optimal Control, Market Simulation.
1 Introduction
Over the last 10-15 years, algorithmic trading and high-frequency trading (HFT) has become the most common method for completing trade transactions in the worlds leading financial markets. Trading algorithms can place many different types of trade orders, but this paper will solely focus on a strategy which uses market orders (MO) and limit orders (LO) to simplify the strategy simulation process. MOs are trade orders executed at the best available prices and LOs are trade orders which rest in the limit order book (LOB) until cancelled or executed by a corresponding MO. See Figure 1 for a visual description of the LOB in the E-mini S&P 500 contract (ES), where we selected an arbitrary time during the most active US stock market trading hours (9:30 EST - 16:00 EST). We will refer to these futures contracts by their ticker symbols from now on.
Mathematical finance experts have long used stochastic processes to model the variables in algorithmic and HFT markets, where Cartea et al., (2015) brings together a whole range of Stochastic Optimal Control (SOC) examples with many variations of trading problems related to liquidation, acquisition, market making, volume-targeted trading, pairs trading, statistical arbitrage and order imbalance. The type of problem that this paper will focus on is the Market-Making (MM) style problem given in section 10.4.2 of Cartea et al., (2015), where the stochastic control is whether or not the market-maker should be posted on the best bid/ask. This problem was chosen as it highlights two of the major issues in MM that are often overlooked and oversimplified in the literature, the fill probability and the adverse fill. These can be summarized as follows:
  • Fill Probability: Measures the probability of a traders LO being filled when they’re posted in the LOB. This often depends on the depth in the LOB and the average trade size. As can be seen in Figure 1, one measure of the depth is the size of the bid and ask queues displayed in the LOB. One could also look at the volume traded relative to these LO sizes and in Table 1 we show the mean and median trade size per contract in four of the most liquid US futures contract, where we used LOB data from April 9th 2024 to calculate this.
    Contract Mean Trade Size Median Trade Size
    ES 3.47 1
    NQ 1.54 1
    CL 1.70 1
    ZN 16.01 2
    Table 1: The mean and median trade sizes in four of the most liquid Chicago Mercantile Exchange (CME) futures contracts.
  • Adverse Fills: A form of adverse selection which occurs when a passive MMs LO is picked off and they receive disadvantageous prices on their filled trade orders. In other words, immediately after their trade order is filled, their new trade position is out of the money when marked-to-market. We will show how the adverse fill occurs on the majority of trade order fills across many different assets, and based on this empirical evidence, we devised a method for including these adverse fills in our proposed trading simulation environment.
Figure 1: A snapshot of the LOB data on April 24th, 2024 for the ES futures contract which expires in June 2024, where the x-axis shows the size of the LOs and the y-axis the price.
This paper will proceed as follows. We begin by discussing the significance of the paper along with a motivating example. In section 2, we will go over some of the literature on fill probabilities and adverse fills. In section 3, we introduce a SOC MM problem, which we will use for our simulation analysis. In section 4, we explain how one can create a simulation environment by discretizing the continuous processes given in section 3. Here we compare the simulation environment from Cartea et al., (2015) and Jaimungal, (2019), which we will refer to as the benchmark environment, with our improved simulation environment. Then, in section 5, we show the results of the performance of the trading strategy in both simulation environments. Lastly, we finish this paper off with some concluding remarks and ideas for future research.
1.1 Motivation and Significance of the Paper
The goal of this paper will be to contribute to the existing literature on the trading strategy simulation process by portraying a more accurate way to simulate performance. Most trading simulation frameworks in the literature model price processes and market orders independently, which can often lead to significantly inflated performance measures. The first problem here is that adverse fills often end up being excluded from the trading strategy simulation process, and, secondly, unrealistic fill probabilities are often used for every other type of trade order fill, which we will refer to as non-adverse fills. To stress the importance of tracking adverse fills in the simulation process, one must recall that in a real live trading environment, if a LO is posted at a specific price, and the price of the asset ends up going through this price, you are essentially guaranteed to have gotten a fill at a worse price than the current price. This is merely based on the rules of the LOB system. Thus, excluding any trade order fills of this nature, from the trading strategy simulation process, would create an unrealistic fill tracking process.
To go into a bit more detail on the importance of this, we give an example by analysing some limit order fill data from a professional simulator environment, using the Trading Technologies (TT) software platform. The TT Simulation environment is a SaaS technology trading platform developed by a company called Trading Technologies. As well as enabling users to trade in live markets, TT has developed its own simulation environment where users can test strategies without risking real capital. Simulation performance results here are often highly in-line with what would happen in reality, although it is important to point out that a major exception would be for strategies that aim to trade high relative volume at specific prices. Using the TT platform, we will give some empirical evidence on how often adverse and non-adverse fills occur based on a sample of simulated trades throughout a given trading day. TT can quite accurately simulate when you are likely to receive a LO fill. Put simply, whenever a LO is placed, it is placed in the LOB and given an imaginary queue position which would be the same as what one would have been given in reality. Then, a LO is considered filled when the trade order behind the imaginary order in the queue is filled in reality. This method can simulate fills quite realistically, particularly if the order size is relatively small, as this traders flow is unlikely to have effected prices in reality.
To show the likelihood of adverse fills occurring, we start by simulating a simple trading strategy in ES, NQ, CL and ZN, which posts limit orders close to the best bid/ask over a single trading day. For ES and CL, the strategy placed static orders on the bid and ask 4 ticks away from each other, in NQ 16 ticks away, and in ZN 1 tick away, while also placing orders at the same filled price again once a fill occurs at a different price. We picked these levels so that the basic posting strategy would trade low relative volume throughout the day, but at the same time, complete enough trades so that we can study the empirical limit order fill data. Each fill that precedes a disadvantageous move in the price will be considered adverse, while every other fill will be considered non-adverse. More specifically, when a LO is filled on the bid (ask) and the next new bid (ask) price is lower (higher), we consider this fill to be adverse, as the first change in the marked-to-market value of the position is negative. By marked-to-market, we used the price that the position can immediately be liquidated at using a MO, which involves crossing the spread. In order to check this, we iterated over each future LOB event until the first change in the marked-to-market value for each individual trade was nonzero. See Table 2 for the results, where we can see that a significant portion of the total number of LO fills in ES, NQ, CL and ZN were adverse.
Trade Order Fills
Date Contract Total number of LO fills Number of adverse fills Number of non-adverse fills
2024/04/24 ES Jun24 941 767 174
2024/04/25 NQ Jun24 1929 1269 660
2024/04/23 CL Jun24 625 518 107
2024/04/24 ZN Jun24 224 199 25
Table 2: The different types of LO fills for the basic posting strategy over active trading trading hours (8:30 EST - 16:00 EST).
Thus, based on the above empirical results, it is quite obvious that adverse fills could significantly effect the performance of any short-term intra-day trading strategy posting LOs, which leads us to believe that a trading simulation environment should account for this. Later on, we develop an improved simulation environment in a MM setting, where we show that the results significantly change when the simulation process tracks adverse fills, as well as utilizing non-adverse fill probabilities. We chose to perform our analysis in a MM setting, since MM strategies often trade passively, in the sense that MMs often post a large amount LO’s in the LOB and aim to trade a large amount volume throughout any given day.
2 Literature Review
There are some recent papers in optimal MM that try to incorporate the effects of fill probabilities and adverse selection into the LOB system and the trading strategy process, while there also some that use SOC theory to setup the trading problem. For example, Bulthuis et al., (2017) devises a model for analyzing LOs and MOs and how they create a permanent and temporary price impact, while also including a component for unfilled LOs, under a SOC framework. Cartea et al., 2018a developed a volume imbalance indicator to make predictions about future MOs. They designed a Markovian pure jump model of the price process, under a SOC framework, including variables such as the spread, LO and MO arrivals and volume imbalance. Roldan Contreras and Swishchuk, (2022) study various types of SOC trading problems, including MM, where they aim to portray some of the unique characteristics on how asset prices move through a dependence on recent activity and jump features. Their optimal solutions are given in terms of the arrival rates of trade orders, which they model utilizing the General Compound Hawkes Process framework. Cartea et al., 2018b , Cartea et al., (2015) and Cartea et al., (2014) develop models which assumes HFT has superior information where they can utilize this to post and cancel LOs at a quicker rate. Their models analyze factors such as arriving MOs, trade clustering and variables that can lead to a change in the shape of the LOB. They create a short-term alpha predictor with the hope of predicting future adverse fills so that they can avoid them. While collecting superior information like this may help in certain instances when making future price predictions, it cannot avoid regular random day-to-day trade order flow where it is normally too late to change the order posting strategy when random spikes in activity instantly enter the market. Predictions, in general, can often be wrong too. The instant nature of trading activity means it is often too late to change the LO posting strategy when random spikes occur. We will discuss this in more detail later from section 4 onward.
From a non SOC standpoint, there is also quite a bit of research on LOBs and fill probabilities. For a broad survey on how LOBs are modelled, see Gould et al., (2013). They discuss a large portion of the different methods used to model LOBs from a wide array of empirical and theoretical research studies. They also highlight the many shortcomings and obvious empirical facts that are yet to be included within these models. Cont and De Larrard, (2013) devised a stochastic framework for modelling the dynamics of the LOB where they made computations based on the arrival of MOs and LOs through a Markovian queueing system. Their setting made it easy to compute numerous important variables such as the distribution for the time length between changes in price, the distribution and autocorrelation for the changes in price and predictions for midprice increases/decreases. In Swishchuk and Vadori, (2017) they extend their framework by creating an arbitrary distribution for the inter-arrival times of LOB events, as well as creating a dependence structure between new and previous LOB book events, under a Markov Renewal process framework. In Swishchuk et al., (2019), they introduce two types of Hawkes processes: compound and regime-switching. Through these models they express price volatility in terms of parameters describing arrival rates and price changes. Maglaras et al., (2022) uses a recurrent neural network to predict the fill probabilities of LOs. It is the first study that directly predicts LO executions via the distribution of fill times. Arroyo et al., (2024) develops a model using survival analysis which maps time-varying features of the LOB to the distribution of fill times of limit orders. Their method uses a convolutional-transformer encoder and a monotonic neural network decoder. Lokin and Yu, (2024) proposes a queueing system method for calculating the fill probabilites of LOs posted at different levels in the LOB. Their probability model covers a wide array of scenarios including midprice changes, probabilities for orders posted at the best bid/ask and at deeper levels in the LOB. Hoffmann, (2014) devises a framework whereby you have fast and slow traders interacting together, where the fast traders can quickly revise their quotes after new information enters the market, which thus creates the ability to avoid being picked off. Cont et al., (2014) examines the impact of LOB events and that price movements occur mostly due to order flow imbalance. They portray how this could be used as a measure of adverse selection in LOs. Easley et al., (2012) discusses a method for measuring how toxic order flow can effect the MM adversely, based on volume imbalance and intensity of orders. They argue it is a good model for predicting toxic order flow over a short-term horizon. Brogaard et al., (2019) devises a model for analyzing how MOs and LOs entered by HFT and non-HFT traders induce price discovery, which they argue mostly comes from LOs, and the majority of these orders come from HFTs. They portray how a decrease in the volume of LOs entering the market often leads to increased volatility in the price.
As described above, some forms of adverse selection in relation to adverse fills have been incorporated into different frameworks in the literature. However, in terms of correcting the trading process or the LOB models for adverse fills, we have seen no prior work. No matter how good a proposed trading model is, it will be virtually impossible to avoid all adverse fills. DeLise, (2024) is the first paper to give an in-depth analysis of the likelihood of adverse fills, where they show some empirical evidence for how often adverse fills occur. They portray how the majority of trade order fills from the financial asset they study, ZN, are, in fact, adverse. We extended their ideas by also analysing the rate of adverse fills in ZN in Table 2, as well as some additional products, to further justify the empirical evidence on adverse fills, which made it clear it’s paramount to include these in any short-term trading strategy evaluation process.
3 A Stochastic Optimal Control Market-Making Problem
A MM problem, in the SOC sense, involves a financial market player who would like to maximize their terminal wealth by trading a large amount of a financial asset, often using many LOs. This type of trader is often referred to as a liquidity provider, which refers to their ability to post a large amount of LOs throughout the LOB. In the example problem we give, the control variable is whether to be posted on the best bid/ask, which is often a decision a MM tries to make. What makes this particular trading problem special is that it begins to incorporate adverse selection in the trading process, which is a big hurdle for a market-maker in reality.
Firstly, we would like to introduce the MM problem we studied, as given in Cartea et al., (2015). We begin by describing the key stochastic processes as follows,
  • $\delta^{\pm}=(\delta_{t}^{\pm})_{\{0\leq t\leq T\}}$, where $\delta_{t}^{\pm}\in\{0,1\}$, is the agents control process, which tells the market-maker whether or not to be posted at the best bid/ask.
  • $Q^{\delta}=(Q_{t}^{\delta})_{\{0\leq t\leq T\}}$ is the agents’ controlled inventory process and is impacted by how much the agent trades.
  • $S^{\delta}=(S_{t}^{\delta})_{\{0\leq t\leq T\}}$ is the controlled midprice process of the financial asset being traded, which is also impacted by how much the agent trades.
  • $M^{\pm}=(M_{t}^{\pm})_{\{0\leq t\leq T\}}$ is the MO arrival process. Here, buy (+)(+)( + ) and sell ()(-)( - ) MOs arrive at Poisson times with intensities $\lambda^{\pm}$. The MO arrival process is essentially a counting process for all the arriving MOs.
  • $N^{\delta,\pm}=(N_{t}^{\delta,\pm})_{\{0\leq t\leq T\}}$ is the trade order fill process, which is also a counting process but with a dependence on the LO posting control. Here, sell (+)(+)( + ) and buy ()(-)( - ) LOs are filled when they match with an incoming MO, $M^{\pm}$.
  • $C^{\delta}=(C_{t}^{\delta})_{\{0\leq t\leq T\}}$ is the agents cash process resulting from the execution strategy.
Next, we see how these processes satisfy certain SDEs as explained in Cartea et al., (2015) and Cartea et al., 2018b as follows:
  • The midprice process satisfies the SDE,
    $dS_{t}=(\nu+\alpha_{t})dt+\sigma dW_{t},$ (1)
    Here, the drift is given by a long-term component $\nu$ and by a short-term component $\alpha_{t}$, which is a predictable zero-mean reverting process, and $W_{t}$ is a standard arithmetic Brownian motion.
  • The short-term drift component in equation (1), $\alpha_{t}$, can be specified in many ways and we follow the method in Cartea et al., (2015) and Cartea et al., 2018b for simplicity. Here, it’s assumed that the short-term alpha component is influenced by order flow, where it predicts the price to jump in the direction of arriving market orders. Thus, when buy MOs arrive, it creates an upward jump in the price, and when sell MOs arrive, it creates a downward jump in the price. And so, they model $\alpha_{t}$ as a zero mean-reverting process, with jumps occurring at the arrival times of MOs, where the size of these jumps is random. Thus, in mathematical terms, $\alpha_{t}$ satisfies the SDE,
    $d\alpha_{t}=-\zeta\alpha_{t}dt+\eta dW_{t}^{\alpha}+\epsilon_{1+M_{t}^{-}}^{+}dM_{t}^{+}-\epsilon_{1+M_{t}^{-}}^{-}dM_{t}^{-}.$ (2)
    Here, $\zeta$ and $\eta$ are positive constants, $W_{t}^{\alpha}$ is a Brownian motion independent of all other processes and ${\epsilon_{1}^{\pm},\epsilon_{2}^{\pm},...}$ are i.i.d random variables also independent of all other processes representing the size of these MOs which arrive at an independent constant rate $\lambda^{\pm}$.
  • The controlled inventory process, which changes every time the agents posted LO is filled by an incoming MO, can be stated as follows,
    $Q_{t}^{\delta}=N_{t}^{\delta,-}-N_{t}^{\delta,+},$ (3)
    where $N_{t}^{\delta,-}$ ($N_{t}^{\delta,+}$) represents buy (sell) LO fills that increases (decreases) the agents inventory.
  • The agents cash process satisfies the SDE,
    $dC_{t}^{\delta}=\displaystyle{\left(S_{t}+\frac{\Delta}{2}\right)}dN_{t}^{\delta,+}-\displaystyle{\left(S_{t}-\frac{\Delta}{2}\right)}dN_{t}^{\delta,-}$ (4)
    Here, $\Delta$ is the spread between the best bid and ask, and, as before, $N_{t}^{\delta,\pm}$ represents the counting process for filled LOs.
The agents’ performance criteria, as in Cartea et al., (2015) and Cartea et al., 2018b , is then defined as follows,
[…]
$H(t,c,S,\alpha,q)=\sup_{\delta\in\mathscr{A}}H^{\delta}(t,c,S,\alpha,q),$ (6)
where $\mathscr{A}$ is the set of admissible strategies in which $\delta>0$ and uniformly bounded from above. It is also important to note that there is an inventory constraint in this problem, whereby the inventory process has an upper $\overline{q}$ and lower $\underline{q}$ bound. This essentially sets the maximum position the market maker can take, long or short. Then, by applying the Dynamic Programming Principle (DPP), as explained in Cartea et al., (2015), the value function should satisfy the following Dynamic Programming Equation (DPE),
$\begin{split}0=\left(\partial_{t}+\alpha\partial_{s}+\frac{1}{2}\sigma^{2}\partial_{ss}-\zeta\alpha\partial_{\alpha}+\frac{1}{2}\eta^{2}\partial_{\alpha\alpha}\right)H-\phi q^{2}\\ +\lambda^{+}\max_{\delta^{+}\in\{0,1\}}\{\mathbb{I}_{q>\underline{q}}\mathbb{E}\left[H(t,c+\left(S+\frac{\Delta}{2}\delta^{+}\right)\delta^{+},S,\alpha+\epsilon^{+},q-\delta^{+})-H\right]\\ +(1-\delta^{+}\mathbb{I}_{q>\underline{q}})\mathbb{E}\left[H(t,c,S,\alpha+\epsilon^{+},q)-H\right]\}\\ +\lambda^{-}\max_{\delta^{-}\in\{0,1\}}\{\mathbb{I}_{q<\overline{q}}\mathbb{E}\left[H(t,c-\left(S-\frac{\Delta}{2}\delta^{-}\right)\delta^{-},S,\alpha-\epsilon^{-},q+\delta^{-})-H\right]\\ +(1-\delta^{-}\mathbb{I}_{q<\overline{q}})\mathbb{E}[H(t,c,S,\alpha-\epsilon^{-},q)-H]\}\end{split}$ (7)
subject to the terminal and boundary conditions:
$H(T,c,S,\alpha,q)=c+q\displaystyle\left(S-\left(\frac{\Delta}{2}+\varphi q\right)\right).$ (8)
Note that the expectations are over the i.i.d. random jump sizes ϵ±superscriptitalic-$\epsilon^{\pm}$. Each line in the DPE can be explained as follows:
  • Line 1: Characterizes the drift and diffusive parts of the midprice and short-term alpha processes, along alpha’s mean-reverting component. The last term on this line includes the running inventory penalty component, which only effects the inventory process.
  • Line 2: Characterizes the change in the value function if the agent has a sell LO posted and the effect of an arriving buy MO filling the agent’s posted sell LO. This filled LO creates a jump in the short-term alpha component. The maximisation term characterizes the agent’s control, which determines whether to post a LO at the best offer.
  • Line 3: Characterizes the change in the value function when an MO arrives but the agent is not posted, in which case only the short-term alpha jumps.
  • Line 4-5: Analogous to lines 2-3 but for the buy side of the LOB.
In Cartea et al., (2015), they use an ansatz to simplify the DPE in equation (7). This ansatz splits out the accumulated cash, the book value of the shares which are marked-to-market at the midprice and the added value from making markets optimally as follows:
$H(t,c,S,\alpha,q)=c+qS+h(t,\alpha,q)$ (9)
Substituting this into the DPE in equation (7) gives us,
$\begin{split}0=\displaystyle\left(\partial_{t}-\zeta\alpha\partial_{\alpha}+\frac{1}{2}\eta^{2}\partial_{\alpha\alpha}\right)h+\alpha q-\phi q^{2}\\ +\lambda^{+}\max_{\delta^{+}\in\{0,1\}}\displaystyle\left\{\mathbb{I}_{q>\underline{q}}\mathbb{E}\displaystyle\left[\delta^{+}\frac{\Delta}{2}+h(t,\alpha+\epsilon^{+},q-\delta^{+})-h(t,\alpha+\epsilon^{+},q)\right]\right\}\\ +\lambda^{-}\max_{\delta^{-}\in\{0,1\}}\displaystyle\left\{\mathbb{I}_{q<\overline{q}}\mathbb{E}\displaystyle\left[\delta^{-}\frac{\Delta}{2}+h(t,\alpha-\epsilon^{-},q+\delta^{-})-h(t,\alpha-\epsilon^{-},q)\right]\right\}\\ +\lambda^{+}\mathbb{E}[h(t,\alpha+\epsilon^{+},q)-h(t,\alpha,q)]\\ +\lambda^{-}\mathbb{E}[h(t,\alpha-\epsilon^{-},q)-h(t,\alpha,q)]\end{split}$ (10)
subject to the terminal condition,
$h(T,\alpha,q)=-q\displaystyle\left(\frac{\Delta}{2}+\varphi q\right)$ (11)
Lastly, by letting $\delta=0$, the optimal postings can be represented in a compact form due to the structure of the maximization terms as follows,
$\begin{split}\delta^{+,*}(t,\alpha,q)=\mathbb{I}_{\{\frac{\Delta}{2}+\mathbb{E}[h(t,\alpha+\epsilon^{+},q-1)-h(t,\alpha+\epsilon^{+},q)]>0\}\cap\{q>\underline{q}\}}\\ \delta^{-,*}(t,\alpha,q)=\mathbb{I}_{\{\frac{\Delta}{2}+\mathbb{E}[h(t,\alpha-\epsilon^{-},q+1)-h(t,\alpha-\epsilon^{-},q)]>0\}\cap\{q<\overline{q}\}}\\ \end{split}$ (12)
Here, we can see that the decision to post depends on the half-spread plus the expectation over the value functions computed at $\alpha\pm\epsilon^{\pm}$, where an arriving MO leads to a positive/negative jump in $\alpha$.
4 Trading Simulation Environment
In this section we discuss the trading simulation environment used to test the performance of the optimal MM strategy. In Cartea et al., (2015) and Jaimungal, (2019), they create a simulation environment in order to back-test this strategy, which we used as our benchmark. We will make one update to the benchmark by using real LOB data, rather than simulated prices. This enabled us to back-test the strategy on real financial futures contracts data later on. We will discuss the extensions we made in our improved simulation environment, which we believe more accurately portrays how this type of MM strategy would perform in reality, albeit, still with some limitations. Specifically, we will describe how our framework tracks the adverse fills, as well as an improved fill probability structure related to non-adverse fills. We will then proceed to show how this can have a major effect on the performance of the trading strategy.
First, we briefly describe the simulation environment used in Cartea et al., (2015) and Jaimungal, (2019) for the optimal MM problem described in section 3, along with the extensions we made in a little more depth. In order to simulate the performance of the MM strategy, one must first discretize the continuous-time setup given in section 3. To begin, the main components of the simulation environment are as follows,
  • Asset price: For this purpose, we will use Futures LOB data, which was collected at the proprietary trading firm Futures First Inc. The products we will focus on include the near-term expiry contracts for ES, NQ, CL and ZN. Market makers often tend to focus on the more liquid products like these, in particular for HFT, as it will enable them to trade more volume. We resampled our price data into a 1 second time interval, which was then subsequently used as the time step size in our discrete environment.
  • LO postings: Here the decision/control variable is stored, which tells us whether to be posted at the best bid/ask at any given time step. The formula for the optimal control is given in equation (12), where the values come from the numerical solution to equation (10). Since a numerical solution is discrete, the values from this solution can be used directly to determine whether to be posted or not. See Jaimungal, (2019) for how this problem is solved numerically.
  • Trade order fills: The benchmark trading simulation environment essentially assumes all fills are non-adverse, unless in the unlikely event a fill occurs at the exact same time as an adverse price movement. In our improved environment, however, we include adverse fills, as well as non-adverse fills, in the fill tracking process. The benchmark environment also assumes that all trade orders go to the front of the LOB queue, thus, their LOs will be the first to be filled. In most liquid financial assets, this is not very likely to occur. Based on this and using the empirical evidence from Tables 1 and 2 earlier, we will attach a binomial probability $p$ to all non-adverse fills. And so, non-adverse fills can be defined as follows,
    $NFA_{t}=\sum_{i=1}^{N}\delta^{+}_{t_{i}}\mathbb{I}_{\{M^{+}_{t_{i}}=1\}}*p$ (13)
    and
    $NFB_{t}=\sum_{i=1}^{N}\delta^{-}_{t_{i}}\mathbb{I}_{\{M^{-}_{t_{i}}=1\}}*p,$ (14)
    for $i=0,...,N$, where $N$ is the total number of time steps. Here, $NFA_{t}$ and $NFB_{t}$ are the counting processes for all non-adverse fills on the best ask and best bid, respectively. $p$ can be defined as:
    $p=P(N_{t_{i}}^{\delta,\pm}|\delta_{t_{i}}^{+}\mathbb{I}_{\{AS(t_{i})\geq AS(t_{i+1})\}}=0,\delta_{t_{i}}^{-}\mathbb{I}_{\{BS(t_{i})\leq BS(t_{i+1})\}}=0)$ (15)
    which is the probability of getting filled given the fill is non adverse. In Cartea et al., (2015) and Jaimungal, (2019), although not explicitly defined, this probability would be one, i.e., $p=1$, and thus we keep it like this in the benchmark environment. Next, we will keep track of the adverse fills in the following manner:
    $AFA_{t}=\sum_{i=1}^{N}\delta_{t_{i}}^{+}\mathbb{I}_{\{AS(t_{i})<AS(t_{i+1})\}},$ (16)
    and
    $AFB_{t}=\sum_{i=1}^{N}\delta_{t_{i}}^{-}\mathbb{I}_{\{BS(t_{i})>BS(t_{i+1})\}},$ (17)
    for $i=0,...,N$, where $AFA_{t}$ and $AFB_{t}$ are counting processes for adverse fills on the ask and adverse fills on the bid, respectively. Here, $BS_{t_{i}}$ and $AS_{t_{i}}$ represent the bid and ask prices, respectively. To understand this intuitively, at each point in time, the trading strategy has a LO posted on the best bid (ask) $\delta_{t_{i}}^{-}(\delta_{t_{i}}^{+})$ and the asset price at time step $t_{i+1}$ is lower (higher) than at time $t_{i}$, the traders LO was then filled at the price $BS_{t_{i}}(AS_{t_{i}})$. This is quite logical because, in reality, based on the rules of the LOB system, the asset price can not move below (above) a traders posted bid (ask) LO without filling their LO first. All other fills the trader receives will be considered non-adverse fills. Finally, in order to determine whether a fill occurred or not, we combine the two fill tracking formulas as follows,
    $N_{t}^{\delta,+}=\sum_{i=1}^{N}max(AFA_{t_{i}},NFA_{t_{i}})$ (18)
    and
    $N_{t}^{\delta,-}=\sum_{i=1}^{N}max(AFB_{t_{i}},NFB_{t_{i}})$ (19)
    where $N_{t}^{\delta,+}$ and $N_{t}^{\delta,-}$ are now counting processes collecting all the trade order fills that occur throughout this particular trading strategy.
For the processes involving MO arrivals, inventory, short-term alpha and cash, we follow the method given in Jaimungal, (2019), where they essentially discretize the equations (2)-(4), as well as devising a unique way to simulate MOs.
5 Results
Here, we will discuss the results from testing the SOC MM strategy in our improved simulation environment and we will compare this to the benchmark environment. First, we would like to briefly summarize the numerical solution under our parameter set and we provide a visual in Figure 2, which looks similar to the solution given in Cartea et al., (2015) and Jaimungal, (2019). One can see how the optimal solution varies as a function of time ($t$), short-term alpha ($\alpha$) and inventory($q$). The left (right) Figure portrays a visual of LOs posted on the best offer (bid), where a sell (buy) LO is posted if the MMs inventory is above (below) the surface. The images look almost identical as the arrival rates of MOs are the same on both sides of the LOB. Far from maturity, the strategy is mostly independent of time, but one can see that as the MM approaches the maturity time ($T$), the optimal strategy is essentially independent of the short-term alpha ($\alpha$), since the MM is solely focused on liquidating any remaining position.
Figure 2: Optimal solution as a function of time, short-term alpha and inventory.
As discussed earlier, the main difference in our improved simulation environment is that adverse fills are now included in the strategy simulation process, determined based on equations (16)-(17). Our aim in this section is to show how adverse fills and fill probabilities on non-adverse fills can significantly effect the performance of the trading strategy. We will discuss the results for CL in depth here, where the same repeated analysis for ES, NQ and ZN can be found in the appendix. Based on the empirical evidence for the distribution between adverse and non-adverse fills given in Table 2, we also set the non-adverse fill probability to $p=0.2$. See Table 3, below, for a list of the parameter values used. We kept most of the parameters similar to Cartea et al., (2015) and Jaimungal, (2019), however, we decreased the time step and increased the maturity time, in order to make the length of time more realistic to the data. We also decreased the intensity of the MOs as another way to bring the distribution of adverse and non-adverse fills a bit closer to reality. A few slight adjustments to some of the other parameters were also needed to keep the numerical solution stable.
Parameters
Parameter Value Parameter Value
$\sigma$ 0.005 T 120 seconds
POV 0.2 Nq 7
$\zeta$ 0.05 $\epsilon$ 0.002
$\eta$ 0.001 $\Delta$ 0.01
$\varphi$ 0.01 $\phi$ 0
$\lambda^{+}$ 0.1458 $\lambda^{-}$ 0.1458
Ndt 120 dt 1 second
Table 3: Simulation parameters.
Now, we will review the strategy simulation from our improved simulation environment alongside the performance of the benchmark environment. In Figure 3, we begin by showing a snapshot of the strategy over a random 120 second path in CL, where one can see when the market-maker is posted on the best bid/ask and when they receive trade order fills. The left and right Figures show the strategy simulation in the benchmark and improved environments, respectively. Here, the green (blue) lines indicate when the market-maker is posted on the best bid (ask), the filled circles indicate when the market-makers LO is filled on the best bid/ask, and the unfilled circles indicate times the market-maker would have been filled if they had a LO posted. One noticeable feature, in the left Figure, is that whenever the agent is posted and price moves through their order, they do not automatically receive a fill. This, as we mentioned earlier, is because MOs are simulated independently from the price process in Cartea et al., (2015), Cartea et al., 2018b and Jaimungal, (2019), and is contrary to what would happen in reality. And so, all the fills in this left Figure would be referred to as non-adverse fills. In the right Figure, one can see where all the adverse fills would have occurred (denoted by AFB and AFA). Note, this can also alter the posting strategy later on because the inventory process now evolves differently to how it would in the benchmark environment.
Figure 3: A snapshot of a random path of the strategy in the benchmark (left) and improved (right) simulation environments. The top line (black/blue) indicates the best ask, the middle line (black) the midprice and the bottom line (black/green) the best bid. When the top (bottom) line is blue (green), the agent is posted on the best ask (bid). Closed circles indicate fills received when the market-maker is posted and open circles indicate fills that the market-maker would’ve received if they were posted.
It’s quite apparent that the timing of the trade order fills has changed and many trade order fills occur at very unfavourable prices. Earlier, in our simple trading example, where we gave results in Table 2, we saw there that a overwhelming majority of the fills for the strategy simulation in CL were adverse, as well as for ES, NQ and ZN. This test was also conducted on the exact same trading day as this simulation. In Table 4, we show the number of each type of fill that occurred within our improved simulation environment. We can see that using a much lower non-adverse fill probability of $p=0.2$ created a similar distribution of fills between non-adverse fills and adverse fills as in Table 2, and we will soon see that the overall performance has now dropped significantly. This trend also persists in the strategy simulations using ES, NQ and ZN data, as can be seen in the appendix. Based on our empirical evidence, we believe these results more accurately portray how the strategy would perform in reality. However, we may not want to be overly pessimistic by thinking the SOC MM strategy should have the same number of adverse fills as the basic posting strategy, because the SOC MM model has a short-term alpha predictor which could potentially predict some of these adverse fills. However, it’s extremely unlikely that any short-term alpha predictor could foresee all future adverse fills.
Fill Type Amount
AFA 3950
NFA 825
AFB 3943
NFB 813
Table 4: Number of trade order fills that were adverse and non-adverse.
Next, in Figure 4, one can see the wealth and inventory path for the same random 120 second run of the strategy. Here, again, the left and right figures are under the benchmark and improved simulation environments, respectively. The wealth path is marked-to-market at each time step as in equation (4) and the inventory path changes whenever a trade order fill is received as in equation (3). Here, the market-makers wealth path in the left Figure finishes positive, and in the right Figure it finishes negative, over the same 120 second path. The inventory paths’ show that the benchmark environment traded a lot less than our improved environment, which is mostly due to the additional adverse fills this trader received. The additional adverse fills also significantly outweighed the reduced number of non-adverse fills. And so, it’s quite apparent that these paths now evolve very differently. Lower non-adverse fill probabilities and adverse fills has significantly effected how these processes evolved.
In order to tell the full story, we show a histogram, in Figure 5, of the terminal cash values from all 330 simulation runs of the strategy over this trading day. Each bin indicates how often the strategy paths attained a certain P&L. Here, again, the left and right Figures show the benchmark and improved simulation environments, respectively. One can see that, overall, the strategy performed reasonably well in the benchmark environment and much worse in our improved environment. We believe the performance metric in the benchmark environment is over-inflated, as the market-maker received a lot more non-adverse fills than likely in reality, as well not receiving any adverse fills, which clearly significantly effects the performance. In the right Figure, we can see that the strategy does not perform as well and regularly suffers large losses, relatively speaking. This confirms to us that simulating the performance of a short-term style trading strategy, in particular one involving posting many LOs, is very misleading if it does not track adverse fills and include more accurate non-adverse fill probabilities. This is just as evident in ES, NQ and ZN, where similar results for these assets can be seen in the appendix.
Figure 4: Wealth and inventory processes for the same random path of the strategy under the benchmark (left) and improved (right) simulation environments, respectively.
Figure 5: Terminal cash histogram for the benchmark (left) and improved (right) simulation environments over all 330 simulation paths.
6 Conclusion and Future Recommendations
In this paper we simulated the performance of a SOC optimal MM strategy that both included and excluded adverse fills, as well as adjusting the probability of a non-adverse fill. Many MM style strategies often trade passively by posting many LOs throughout the LOB, thus the results here are particularly relevant to these types of strategies, but also to any trading strategy that posts a large amount LOs. Models in mathematical finance often make simplifying assumptions in order to find robust solutions to a problem and while many don’t often effect the overall performance too much, some certainly can. In our setting, modelling the price process and the arrival of MOs independently can create large inaccuracies, which is clear from our simulation results. Specifically, our finding shows that excluding adverse fills and non-adverse fill probabilities can significantly inflate the performance of a trading strategy. Adverse trade order fills are never completely avoidable for a trader who devises a LO posting strategy and are particularly apparent in short-term strategies. Although we can try to predict short-term price movements using methods like the short-term alpha predictor given in Cartea et al., (2015) and Cartea et al., 2018b , large orders often randomly enter the market taking out price levels on a whim. Thus, if a MM or trader is posted at these price levels they will receive adverse fills as there usually isn’t enough time to cancel the orders, regardless of some of the speed advantageous some HFTs may have.
In regard to future research recommendations, we think the results could be improved if a more dynamic MM model were used, where the MM can update the important parameters over time. It is clear from our analysis and the literature that LOB data has a clustering effect, thus this could enable the MM to update their non-adverse fill probabilities, as well as improving the short-term alpha predictors ability to predict adverse moves. Papers like Arroyo et al., (2024) and Maglaras et al., (2022) are some new initial works that aim to estimate fill probabilities using deep learning approaches, which seem to develop some promising ideas for devising a more dynamic model. One should also assess the performance of any short-term alpha predictor used to predict adverse moves, as this is one of the key components of the model in this paper. In the SOC framework, this simple short-term alpha predictor helped bring tractable results, but it is unlikely to be that simple to make such price predictions in reality. Lastly, one could also try to improve the price process modelling aspect, by using a jump-diffusion or Levy-style model. These price processes have been proven to more accurately mimic LOB dynamics rather than the general diffusion models, although sometimes more difficult to apply directly in a SOC model.
Acknowledgements
We would like to thank MITACS and NSERC for research funding. We would specifically like to thank Futures First for collaborating in this MITACS project, where we would also like to give a special mention to Timothy DeLise (PhD candidate at the University of Montreal and former MITACS participant in projects at Futures First) and Myles Sjogren (Quantitative Analyst at Futures First) in helping to set-up parts of the coding framework for deploying the basic posting algorithm described in section 1.1.
Declarations of Interest
The authors declare no conflicts of interest.
Appendix A Appendix: ES, NQ and ZN Strategy Simulations
Figure 6: A snapshot of the same random path of the strategy in the benchmark (left) and improved (right) simulation environments. The top line (black/blue) indicates the best ask, the middle line (black) the midprice and the bottom line (black/green) the best bid. When the top (bottom) line is blue (green), the agent is posted on the best ask (bid). Closed circles indicate fills received when the market-maker is posted and open circles indicate fills that the market-maker would’ve received if they were posted.
Figure 7: Wealth process for the same random path of the strategy under the benchmark (left) and improved (right) simulation environments.
Figure 8: Inventory process for the same random path of the strategy under the benchmark (left) and improved (right) simulation environments.
Figure 9: Terminal cash histogram in the benchmark (left) and improved (right) simulation environments for all 330 simulation paths.
Fill Type Amount Fill Type Amount Fill Type Amount
AFA 5904 AFA 12319 AFA 478
NFA 789 NFA 840 NFA 825
AFB 5873 AFB 12370 AFB 468
NFB 774 NFB 781 NFB 821
Table 5: Number of fills received over each 120 second run of the strategy that were adverse and non-adverse in ES, NQ and ZN, from left to right.

### v2

@@
 In this paper, we study the effects of fill probabilities and adverse fills on the trading strategy simulation process.
-We specifically focus on a stochastic optimal control market-making problem and test the strategy on ES (E-mini S&P 500), NQ (E-mini Nasdaq 100), CL (Crude Oil) and ZN (10-Year Treasury Note), which are some of the most liquid futures contract listed on the CME (Chicago Mercantile Exchange).
-We provide empirical evidence which shows how fill probabilities and adverse fills can significantly effect performance, and propose a more prudent simulation framework for dealing with this.
-Many previous works aim to measure different types of adverse selection in the limit order book, however, they often simulate price processes and market orders independently.
+We specifically focus on a stochastic optimal control market-making problem and test the strategy on ES (E-mini S&P 500), NQ (E-mini Nasdaq 100), CL (Crude Oil) and ZN (10-Year Treasury Note), which are some of the most liquid futures contracts listed on the CME (Chicago Mercantile Exchange).
+We provide empirical evidence that shows how fill probabilities and adverse fills can significantly affect performance and propose a more prudent simulation framework to deal with this.
+Many previous works aim to measure different types of adverse selection in the limit order book (LOB), however, they often simulate price processes and market orders independently.
 This has the ability to largely inflate the performance of a short-term style trading strategy.
-Our studies show that using more realistic fill probabilities, and tracking adverse fills, in the strategy simulation process, more accurately portrays how these types of trading strategies would perform in reality.
+Our studies show that using more realistic fill probabilities and tracking adverse fills in the strategy simulation process more accurately shows how these types of trading strategies would perform in reality.
 Keywords: Algorithmic and High-Frequency Trading, Market-Making, Adverse Selection, Stochastic Optimal Control, Market Simulation.
@@
 Over the last 10-15 years, algorithmic trading and high-frequency trading (HFT) has become the most common method for completing trade transactions in the worlds leading financial markets.
-Trading algorithms can place many different types of trade orders, but this paper will solely focus on a strategy which uses market orders (MO) and limit orders (LO) to simplify the strategy simulation process.
-MOs are trade orders executed at the best available prices and LOs are trade orders which rest in the limit order book (LOB) until cancelled or executed by a corresponding MO.
-See Figure 1 for a visual description of the LOB in the E-mini S&P 500 contract (ES), where we selected an arbitrary time during the most active US stock market trading hours (9:30 EST - 16:00 EST).
-We will refer to these futures contracts by their ticker symbols from now on.
-Mathematical finance experts have long used stochastic processes to model the variables in algorithmic and HFT markets, where Cartea et al., (2015) brings together a whole range of Stochastic Optimal Control (SOC) examples with many variations of trading problems related to liquidation, acquisition, market making, volume-targeted trading, pairs trading, statistical arbitrage and order imbalance.
-The type of problem that this paper will focus on is the Market-Making (MM) style problem given in section 10.4.2 of Cartea et al., (2015), where the stochastic control is whether or not the market-maker should be posted on the best bid/ask.
-This problem was chosen as it highlights two of the major issues in MM that are often overlooked and oversimplified in the literature, the fill probability and the adverse fill.
+Trading algorithms can place many different types of trade orders, but this paper will focus solely on a strategy that uses market orders (MO) and limit orders (LO) to simplify the strategy simulation process.
+MOs are trade orders executed at the best available prices, and LOs are trade orders that rest in the limit order book (LOB) until canceled or executed by a corresponding MO.
+See Figure 1 for a visual description of the LOB in the E-mini S&P 500 contract (ES), where we selected an arbitrary time during the most active trading hours (9:30 EST - 16:00 EST) of the US stock market.
+We will refer to this futures contract by its ticker symbols from now on.
+Mathematical finance experts have long used stochastic processes to model variables in algorithmic and HFT markets, where Cartea et al., (2015) brings together a wide range of Stochastic Optimal Control (SOC) examples with many variations of trading problems related to liquidation, acquisition, market making, volume-targeted trading, pairs trading, statistical arbitrage, and order imbalance.
+The type of problem that this paper will focus on is the Market Maker (MM) style problem given in Section 10.4.2 of Cartea et al., (2015), where the stochastic control is whether or not the MM should be post one unit (be it shares/lots etc) on the best bid/ask.
+This problem was chosen because it highlights two of the major issues for a MM that are often overlooked and oversimplified in the literature, the fill probability and the adverse fill.
 These can be summarized as follows:
-  • Fill Probability: Measures the probability of a traders LO being filled when they’re posted in the LOB.
-This often depends on the depth in the LOB and the average trade size.
-As can be seen in Figure 1, one measure of the depth is the size of the bid and ask queues displayed in the LOB.
-One could also look at the volume traded relative to these LO sizes and in Table 1 we show the mean and median trade size per contract in four of the most liquid US futures contract, where we used LOB data from April 9th 2024 to calculate this.
+  • Fill Probability: The likelihood that a trader’s LO will be executed after being posted in the LOB.
+This probability is influenced by factors such as the position of the order in the queue, the depth of liquidity at that price level, and the frequency and size of incoming market orders.
+As shown in Figure 1, one way to measure depth is by examining the size of the bid and ask queues in the LOB.
+Another approach is to analyze the ratio of the traded volume to the resting LO sizes.
+Table 1 provides the mean and median trade sizes per contract for four of the most liquid U.S.
+futures contracts, calculated using LOB data from April 9, 2024.
+  • Adverse fills: A form of adverse selection that occurs when a passive MMs LO is executed at a disadvantageous price.
+Specifically, this happens when the MM’s order is “picked off,” which means that immediately after execution, the new trade position is out of the money when marked-to-market.
+We will demonstrate that adverse fills occur in the majority of trade executions across various assets.
+Based on this empirical evidence, we develop a method to incorporate adverse fills into our proposed trading simulation environment, ensuring a more realistic representation of execution risk.
 Contract Mean Trade Size Median Trade Size
@@
 Table 1: The mean and median trade sizes in four of the most liquid Chicago Mercantile Exchange (CME) futures contracts.
-  • Adverse Fills: A form of adverse selection which occurs when a passive MMs LO is picked off and they receive disadvantageous prices on their filled trade orders.
-In other words, immediately after their trade order is filled, their new trade position is out of the money when marked-to-market.
-We will show how the adverse fill occurs on the majority of trade order fills across many different assets, and based on this empirical evidence, we devised a method for including these adverse fills in our proposed trading simulation environment.
 Figure 1: A snapshot of the LOB data on April 24th, 2024 for the ES futures contract which expires in June 2024, where the x-axis shows the size of the LOs and the y-axis the price.
-This paper will proceed as follows.
-We begin by discussing the significance of the paper along with a motivating example.
+The main contributions of the paper can be summarized as follows:
+  • We propose an enhanced simulation framework for evaluating the performance of HFT strategies, particularly in market-making.
+This framework accounts for both adverse and non-adverse trade order fills in a more realistic manner, addressing a key limitation in many existing approaches that assume independence between trade order fills and price processes.
+Our framework highlights how this assumption can significantly affect performance.
+  • We introduce a probability measure for non-adverse trade order fills, refining how these fills are modeled within MM strategies.
+As will be demonstrated in Section 3, simplistic strategies often suffer from excessive adverse fills and are not guaranteed to receive non-adverse fills for every incoming market order.
+While fill probabilities are not new, they should only be applied to non-adverse fills, as adverse fills are guaranteed to occur based on the rules of most LOB systems, a limitation we address here.
+A more realistic probability measure ensures that performance metrics align better with real-world trading conditions.
+  • We analyze a well-studied MM problem using the framework of Cartea et al., (2015), extending their SOC formulation to incorporate non-adverse fill probabilities in the optimal solutions.
+Through an empirical study on four highly liquid CME futures contracts, we demonstrate the significant impact of both non-adverse and adverse fills on performance.
+Although adverse fills could not be directly incorporated into the SOC framework due to necessary dimension reduction techniques, our simulation environment still allows us to assess their implications on performance.
+The rest of this paper will proceed as follows.
 In Section 2, we will go over some of the literature on fill probabilities and adverse fills.
-In section 3, we introduce a SOC MM problem, which we will use for our simulation analysis.
+In Section 3, we begin by discussing the significance of the paper with two simple motivating examples of a MM trading strategy.
+In Section 4, we introduce an SOC MM problem, which we will use for our simulation analysis.
 In Section 5, we explain how one can create a simulation environment by discretizing the continuous processes given in Section 4.
-Here we compare the simulation environment from Cartea et al., (2015) and Jaimungal, (2019), which we will refer to as the benchmark environment, with our improved simulation environment.
-Then, in section 5, we show the results of the performance of the trading strategy in both simulation environments.
-Lastly, we finish this paper off with some concluding remarks and ideas for future research.
-1.1 Motivation and Significance of the Paper
-The goal of this paper will be to contribute to the existing literature on the trading strategy simulation process by portraying a more accurate way to simulate performance.
-Most trading simulation frameworks in the literature model price processes and market orders independently, which can often lead to significantly inflated performance measures.
-The first problem here is that adverse fills often end up being excluded from the trading strategy simulation process, and, secondly, unrealistic fill probabilities are often used for every other type of trade order fill, which we will refer to as non-adverse fills.
-To stress the importance of tracking adverse fills in the simulation process, one must recall that in a real live trading environment, if a LO is posted at a specific price, and the price of the asset ends up going through this price, you are essentially guaranteed to have gotten a fill at a worse price than the current price.
-This is merely based on the rules of the LOB system.
-Thus, excluding any trade order fills of this nature, from the trading strategy simulation process, would create an unrealistic fill tracking process.
-To go into a bit more detail on the importance of this, we give an example by analysing some limit order fill data from a professional simulator environment, using the Trading Technologies (TT) software platform.
-The TT Simulation environment is a SaaS technology trading platform developed by a company called Trading Technologies.
-As well as enabling users to trade in live markets, TT has developed its own simulation environment where users can test strategies without risking real capital.
-Simulation performance results here are often highly in-line with what would happen in reality, although it is important to point out that a major exception would be for strategies that aim to trade high relative volume at specific prices.
-Using the TT platform, we will give some empirical evidence on how often adverse and non-adverse fills occur based on a sample of simulated trades throughout a given trading day.
+Here we compare the simulation environment in Cartea et al., (2015) and Jaimungal, (2019), which we will refer to as the benchmark environment, with our improved simulation environment.
+Then, in Section 6, we show the performance results of the trading strategy in both simulation environments.
+Lastly, this paper ends with some concluding remarks and ideas for future research.
+2 Literature Review
+Recent research on optimal MM problems incorporates fill probabilities and adverse selection in the LOB system and the trading strategy process, with some approaches leveraging SOC theory.
+For example, Bulthuis et al., (2017) develops a model analyzing the price impact of LOs and MOs, incorporating a component for unfilled LOs within an SOC framework.
+Cartea et al., 2018a introduces a volume imbalance indicator to predict future MOs, using a Markovian pure jump model that accounts for the spread, LO/MO arrivals, and volume imbalance.
+Roldan Contreras and Swishchuk, (2022) explores various SOC trading problems, including MM, highlighting how asset prices evolve based on recent activity and jump characteristics.
+Their optimal solutions are expressed in terms of trade order arrival rates, modeled via the General Compound Hawkes Process framework.
+Furthermore, Cartea et al., 2018b , Cartea et al., (2015), and Cartea et al., (2014) examine HFT strategies that aim to leverage superior information to post and cancel LOs more efficiently.
+Their models analyze MO arrivals, trade clustering, and LOB shape dynamics, using a short-term alpha predictor to anticipate and avoid adverse fills.
+However, while such predictive models may offer advantages in some scenarios, they cannot fully account for random, HFT order flow, where sudden spikes in activity occur too quickly for an effective response.
+Predictions can also be unreliable, especially when the short-term alpha predictor lacks a meaningful market edge.
+We explore this issue further in Section 4, where we analyze a similar MM problem.
+Beyond SOC models, there is extensive research on LOBs and fill probabilities.
+For a broad survey, Gould et al., (2013) reviews empirical and theoretical modeling approaches, highlighting key shortcomings.
+Cont and De Larrard, (2013) introduces a stochastic LOB framework using a Markovian queuing system to model MO and LO arrivals, enabling computations of price change distributions, autocorrelations, and midprice movements.
+Swishchuk and Vadori, (2017) extends this by allowing arbitrary inter-arrival distributions and dependencies between LOB events under a Markov Renewal process framework.
+Swishchuk et al., (2019) further enhances this by incorporating compound and regime-switching Hawkes processes, expressing price volatility in terms of arrival rates and price changes.
+Maglaras et al., (2022) employs a recurrent neural network to predict LO fill probabilities, marking the first study to model executions via fill time distributions.
+Arroyo et al., (2024) applies survival analysis with a convolutional transformer encoder and a monotonic neural network decoder to map LOB characteristics to fill time distributions.
+Lokin and Yu, (2024) develops a queuing system for calculating LO fill probabilities at different depths, considering midprice changes and best bid/ask orders.
+Hoffmann, (2014) models interactions between fast and slow traders, where fast traders update quotes rapidly to avoid being picked off.
+Cont et al., (2014) links price movements to order flow imbalance, using it as a measure of adverse selection.
+Easley et al., (2012) proposes a model linking toxic order flow to MM risk via volume imbalance and order intensity, offering a short-term predictive framework.
+Brogaard et al., (2019) analyzes how MO and LO submissions by HFT and non-HFT traders drive price discovery, finding that most price information stems from LOs, particularly from HFTs.
+They also show that reduced LO volume leads to increased price volatility.
+Although some frameworks incorporate adverse fills, no prior work has explicitly adjusted the trading process or LOB models to distinguish between adverse and non-adverse fills.
+Regardless of the quality of the model, it is virtually impossible to completely avoid adverse fills.
+DeLise, (2024) is the first to analyze the likelihood of adverse fills, providing empirical evidence on their frequency.
+They show that in ZN, most trade order fills are adverse.
+We extend their work by examining the adverse fill rate in ZN (Table 2) and additional products, further demonstrating the necessity of incorporating adverse fills in short-term trading strategy evaluations.
+In Section 3, we introduce a method to distinguish between adverse and non-adverse fills through simple MM examples.
+In Section 4, we present a more complex SOC MM problem, and in Section 5, we propose a method to integrate this trade order fill logic into the framework.
+3 Adverse and Non-adverse Trade Order Fills
+Most existing simulation models treat price processes and market orders independently, often inflating performance metrics.
+The primary issue with this is that adverse fills are frequently excluded, while unrealistic fill probabilities are applied to other trade executions, which we refer to as non-adverse fills.
+Accurately tracking adverse fills is crucial because in real trading environments, if a LO is posted at a given price and the asset price moves through it, the order is effectively guaranteed to be filled at a worse price.
+This is a fundamental property of the LOB system.
+Excluding such fills leads to an unrealistic fill-tracking process.
+Similar findings to ours also appear in Law and Viens, (2019) and DeLise, (2024), where the former highlights how diffusion-based market-making models can create “phantom gains”, and the latter provides an empirical analysis of the “negative drift of a limit order fill”, a key factor in an MM’s performance.
+To demonstrate this, consider a real-world LOB observed at 1-second intervals, as we will later in our empirical analysis.
+In most assets, the best bid and ask prices remain unchanged significantly more often than they move at each time step.
+This is supported by empirical analyses in several seminal works, including Biais et al., (1995), Hasbrouck and Saar, (2013), and Cont et al., (2014), among others.
+If market orders are modeled independently of price changes, a trade order fill does not necessarily align with the price moving through an order.
+Unless the model explicitly enforces fills when price levels are crossed, randomly generated executions are unlikely to capture adverse fills, systematically underestimating their frequency.
+This issue worsens as the time step decreases.
+With fewer price changes per step, fills become even less likely to align with price movements, distorting performance assessments, particularly in HFT settings.
+We will also give empirical evidence in Section 5 that demonstrates how failing to account for adverse fills leads to a significant misrepresentation of strategy performance.
+In the remainder of this section, we present two examples to illustrate key concepts and results.
+Section 3.1 introduces a simple example to clarify the distinction between adverse and non-adverse fills, serving purely as an illustration of the two trade-order fill types.
+Section 3.2 explores a basic LO posting strategy for an MM, simulated in a real-time trading environment.
+This second example provides insight into the expected distribution of adverse and non-adverse fills for a simple MM strategy, particularly in the context of relatively small order sizes.
+3.1 Example 1: Simple Market Maker Trade Order Fills
+Consider a simple example that simulates an MM scenario where an LO is always posted on the best bid and ask, and one MO arrives at each time step, randomly filling the MM agent on the bid or ask.
+The best bid and ask prices evolve as a random walk with a fixed 1-cent spread.
+Each fill is classified as adverse (if the next price move is unfavorable) or non-adverse (if the price remains the same or moves favorably).
+It is assumed that the MM agent’s LOs are always at the front of the queue and of size 1, thus, they are fully filled with every incoming MO.
+This can also be interpreted as the MM having a fill probability of 100%.
+See Figure 2, which visualizes bid and ask prices over time, with red markers for adverse fills and green markers for non-adverse fills, clearly illustrating the different trade order fill types.
+In this plot, the best bid and ask prices are represented by blue and yellow lines, respectively.
+Note the following:
+  • The point of this example is merely to highlight the difference between adverse and non-adverse fills.
+  • While the time steps are uniform in this example, they are non-uniform in reality.
+Figure 2: A visualization of the simple MM strategy in Example 1, highlighting the timing of each trade order fill and distinguishing between the different trade order fill types.
+3.2 Example 2: Market Maker Trade Order Fills in a Real-Time Simulation Environment
+To go into a bit more detail on the importance of distinguishing between adverse and non-adverse trade order fills, we give an example by analyzing some LO fill data from a professional simulator environment, using the Trading Technologies (TT) software platform.
+The TT Simulation environment is a trading platform developed by a company called Trading Technologies.
+In addition to providing access to live markets, TT has created a simulation environment that allows users to test strategies without exposing real capital to risk.
+Simulation performance results here are often highly in line with what would happen in reality, although it is important to point out that a major exception would be for strategies that aim to trade high relative volume at specific prices.
+Using the TT platform, we will give some empirical evidence on how often adverse and non-adverse fills occur based on a sample of simulated trades throughout a given trading day (from 8:30 EST-16:00 EST).
 TT can quite accurately simulate when you are likely to receive a LO fill.
-Put simply, whenever a LO is placed, it is placed in the LOB and given an imaginary queue position which would be the same as what one would have been given in reality.
-Then, a LO is considered filled when the trade order behind the imaginary order in the queue is filled in reality.
-This method can simulate fills quite realistically, particularly if the order size is relatively small, as this traders flow is unlikely to have effected prices in reality.
-To show the likelihood of adverse fills occurring, we start by simulating a simple trading strategy in ES, NQ, CL and ZN, which posts limit orders close to the best bid/ask over a single trading day.
-For ES and CL, the strategy placed static orders on the bid and ask 4 ticks away from each other, in NQ 16 ticks away, and in ZN 1 tick away, while also placing orders at the same filled price again once a fill occurs at a different price.
-We picked these levels so that the basic posting strategy would trade low relative volume throughout the day, but at the same time, complete enough trades so that we can study the empirical limit order fill data.
-Each fill that precedes a disadvantageous move in the price will be considered adverse, while every other fill will be considered non-adverse.
-More specifically, when a LO is filled on the bid (ask) and the next new bid (ask) price is lower (higher), we consider this fill to be adverse, as the first change in the marked-to-market value of the position is negative.
-By marked-to-market, we used the price that the position can immediately be liquidated at using a MO, which involves crossing the spread.
-In order to check this, we iterated over each future LOB event until the first change in the marked-to-market value for each individual trade was nonzero.
+Simply put, when an LO is placed, it is assigned an imaginary queue position in the LOB, mirroring the position it would have received in a real market.
+The LO is then considered filled when a real trade order is executed directly behind its imaginary queue position.
+This approach provides a realistic fill simulation, especially for relatively small order sizes, as such trades are unlikely to impact market prices significantly.
+To show the likelihood of adverse fills occurring, we start by simulating a simple trading strategy in ES, NQ, CL and ZN, which posts LOs close to the best bid/ask over a single trading day.
+The main components of the strategy can be described as follows:
+  1. 1.
+    For ES and CL, the strategy placed static orders for 1 lot at bid and ask prices 4 ticks away from each other, in NQ 16 ticks away, and in ZN 1 tick away.
+We picked these levels so that the basic posting strategy would trade low relative volume throughout the day, but at the same time, complete enough trades so that we can study the empirical LO fill data.
+The different tick size spacing in the posted LOs is related to how actively each of these assets trades, as well as how often price changes occur in the asset.
+For example, NQ experiences the most frequent price changes compared to ES, CL, and ZN, while ZN has the fewest and often remains unchanged for extended periods.
+These adjustments allowed us to obtain a reasonable and comparable amount of data for each asset.
+  2. 2.
+    The strategy also reposts orders at the same filled price whenever a fill occurs at a different price.
+A sample set of steps that this strategy could follow, where here the focus will be on CL, can be summarized as follows:
+    1. i.
+      At $t=1$, where $t$ represents a non-uniform timestep determined by event occurrences in the LOB, the MM agent posts two LOs in CL: a buy LO at $81.86$ and a sell LO at $81.90$.
+The price at $t=1$ lies somewhere between these two price levels.
+    2. ii.
+      At $t=2$, the buy LO is filled at $81.86$.
+    3. iii.
+      At $t=3$, the agent places a new buy LO at $81.82$ while keeping the existing sell LO at $81.90$ unchanged.
+    4. iv.
+      At $t=4$, the new buy LO is filled at $81.82$.
+    5. v.
+      At $t=5$, the agent places a new buy LO at $81.78$ and a new sell LO at $81.86$.
+    6. vi.
+      At $t=6$, the new sell LO is filled at $81.86$.
+    7. vii.
+      At $t=7$, the agent places a new buy LO at $81.82$ and, if not already posted, a new sell LO at $81.90$.
+In this example, previously posted orders that are not between the highest posted bid LO and the lowest posted ask LO remain active, though one could also consider canceling resting LOs further from the market.
+    This process continues in a similar fashion throughout the trading day.
+The primary motivation for this LO posting sequence is to prevent the fill data from being skewed by a high concentration of trades occurring at the same price within a short time window.
+This ensures a more representative distribution of adverse and non-adverse fills while also avoiding repetitive oscillatory patterns that may arise over short time intervals.
+As in example 1, each trade order fill that precedes a disadvantageous move in the price will be considered adverse, while every other fill will be considered non-adverse.
+More specifically, when an LO is filled on the bid (ask) and the next new bid (ask) price is lower (higher), we consider this fill to be adverse, as the first change in the marked-to-market value of the position is negative.
 See Table 2 for the results, where we can see that a significant portion of the total number of LO fills in ES, NQ, CL and ZN were adverse.
@@
 Table 2: The different types of LO fills for the basic posting strategy over active trading trading hours (8:30 EST - 16:00 EST).
-Thus, based on the above empirical results, it is quite obvious that adverse fills could significantly effect the performance of any short-term intra-day trading strategy posting LOs, which leads us to believe that a trading simulation environment should account for this.
-Later on, we develop an improved simulation environment in a MM setting, where we show that the results significantly change when the simulation process tracks adverse fills, as well as utilizing non-adverse fill probabilities.
-We chose to perform our analysis in a MM setting, since MM strategies often trade passively, in the sense that MMs often post a large amount LO’s in the LOB and aim to trade a large amount volume throughout any given day.
-2 Literature Review
-There are some recent papers in optimal MM that try to incorporate the effects of fill probabilities and adverse selection into the LOB system and the trading strategy process, while there also some that use SOC theory to setup the trading problem.
-For example, Bulthuis et al., (2017) devises a model for analyzing LOs and MOs and how they create a permanent and temporary price impact, while also including a component for unfilled LOs, under a SOC framework.
-Cartea et al., 2018a developed a volume imbalance indicator to make predictions about future MOs.
-They designed a Markovian pure jump model of the price process, under a SOC framework, including variables such as the spread, LO and MO arrivals and volume imbalance.
-Roldan Contreras and Swishchuk, (2022) study various types of SOC trading problems, including MM, where they aim to portray some of the unique characteristics on how asset prices move through a dependence on recent activity and jump features.
-Their optimal solutions are given in terms of the arrival rates of trade orders, which they model utilizing the General Compound Hawkes Process framework.
-Cartea et al., 2018b , Cartea et al., (2015) and Cartea et al., (2014) develop models which assumes HFT has superior information where they can utilize this to post and cancel LOs at a quicker rate.
-Their models analyze factors such as arriving MOs, trade clustering and variables that can lead to a change in the shape of the LOB.
-They create a short-term alpha predictor with the hope of predicting future adverse fills so that they can avoid them.
-While collecting superior information like this may help in certain instances when making future price predictions, it cannot avoid regular random day-to-day trade order flow where it is normally too late to change the order posting strategy when random spikes in activity instantly enter the market.
-Predictions, in general, can often be wrong too.
-The instant nature of trading activity means it is often too late to change the LO posting strategy when random spikes occur.
-We will discuss this in more detail later from section 4 onward.
-From a non SOC standpoint, there is also quite a bit of research on LOBs and fill probabilities.
-For a broad survey on how LOBs are modelled, see Gould et al., (2013).
-They discuss a large portion of the different methods used to model LOBs from a wide array of empirical and theoretical research studies.
-They also highlight the many shortcomings and obvious empirical facts that are yet to be included within these models.
-Cont and De Larrard, (2013) devised a stochastic framework for modelling the dynamics of the LOB where they made computations based on the arrival of MOs and LOs through a Markovian queueing system.
-Their setting made it easy to compute numerous important variables such as the distribution for the time length between changes in price, the distribution and autocorrelation for the changes in price and predictions for midprice increases/decreases.
-In Swishchuk and Vadori, (2017) they extend their framework by creating an arbitrary distribution for the inter-arrival times of LOB events, as well as creating a dependence structure between new and previous LOB book events, under a Markov Renewal process framework.
-In Swishchuk et al., (2019), they introduce two types of Hawkes processes: compound and regime-switching.
-Through these models they express price volatility in terms of parameters describing arrival rates and price changes.
-Maglaras et al., (2022) uses a recurrent neural network to predict the fill probabilities of LOs.
-It is the first study that directly predicts LO executions via the distribution of fill times.
-Arroyo et al., (2024) develops a model using survival analysis which maps time-varying features of the LOB to the distribution of fill times of limit orders.
-Their method uses a convolutional-transformer encoder and a monotonic neural network decoder.
-Lokin and Yu, (2024) proposes a queueing system method for calculating the fill probabilites of LOs posted at different levels in the LOB.
-Their probability model covers a wide array of scenarios including midprice changes, probabilities for orders posted at the best bid/ask and at deeper levels in the LOB.
-Hoffmann, (2014) devises a framework whereby you have fast and slow traders interacting together, where the fast traders can quickly revise their quotes after new information enters the market, which thus creates the ability to avoid being picked off.
-Cont et al., (2014) examines the impact of LOB events and that price movements occur mostly due to order flow imbalance.
-They portray how this could be used as a measure of adverse selection in LOs.
-Easley et al., (2012) discusses a method for measuring how toxic order flow can effect the MM adversely, based on volume imbalance and intensity of orders.
-They argue it is a good model for predicting toxic order flow over a short-term horizon.
-Brogaard et al., (2019) devises a model for analyzing how MOs and LOs entered by HFT and non-HFT traders induce price discovery, which they argue mostly comes from LOs, and the majority of these orders come from HFTs.
-They portray how a decrease in the volume of LOs entering the market often leads to increased volatility in the price.
-As described above, some forms of adverse selection in relation to adverse fills have been incorporated into different frameworks in the literature.
-However, in terms of correcting the trading process or the LOB models for adverse fills, we have seen no prior work.
-No matter how good a proposed trading model is, it will be virtually impossible to avoid all adverse fills.
-DeLise, (2024) is the first paper to give an in-depth analysis of the likelihood of adverse fills, where they show some empirical evidence for how often adverse fills occur.
-They portray how the majority of trade order fills from the financial asset they study, ZN, are, in fact, adverse.
-We extended their ideas by also analysing the rate of adverse fills in ZN in Table 2, as well as some additional products, to further justify the empirical evidence on adverse fills, which made it clear it’s paramount to include these in any short-term trading strategy evaluation process.
+Thus, the empirical results above clearly demonstrate that adverse fills can significantly impact the performance of any short-term intraday trading strategy that posts LOs.
+This highlights the need for a trading simulation environment that properly accounts for adverse fills, particularly when analyzing a simple MM strategy.
+In Section 4, we introduce an SOC optimal MM problem, where this simple strategy is unlikely to predict future order flow with high accuracy or anticipate price movements as effectively as professional MMs, who typically employ more sophisticated strategies.
+Unfortunately (for academics), most of these strategies remain highly confidential, as revealing them would lead to alpha decay (see Zhou and Lin, (2017) and Sivaramakrishnan et al., (2018) for an empirical analysis of this phenomenon), making them difficult to replicate.
+Although replicating more sophisticated strategies is challenging, their performance assessment measures can still be studied and further improved, as these measures are largely strategy independent.
+Consequently, in our SOC MM scenario, we should expect a similar distribution of adverse and non-adverse fills.
+Later, we develop an improved MM simulation environment, demonstrating that results change significantly when the model explicitly tracks adverse fills and incorporates realistic non-adverse fill probabilities.
+We focus on MM strategies because MMs typically trade passively, posting large volumes of LOs in the LOB and executing a substantial number of trades throughout the day.
 4 A Stochastic Optimal Control Market-Making Problem
-A MM problem, in the SOC sense, involves a financial market player who would like to maximize their terminal wealth by trading a large amount of a financial asset, often using many LOs.
+An MM problem, in the SOC sense, involves a financial market player who would like to maximize their terminal wealth by trading a large amount of a financial asset, often using many LOs.
 This type of trader is often referred to as a liquidity provider, which refers to their ability to post a large amount of LOs throughout the LOB.
 In the example problem we give, the control variable is whether to be posted on the best bid/ask, which is often a decision a MM tries to make.
-What makes this particular trading problem special is that it begins to incorporate adverse selection in the trading process, which is a big hurdle for a market-maker in reality.
-Firstly, we would like to introduce the MM problem we studied, as given in Cartea et al., (2015). We begin by describing the key stochastic processes as follows,
-  • $\delta^{\pm}=(\delta_{t}^{\pm})_{\{0\leq t\leq T\}}$, where $\delta_{t}^{\pm}\in\{0,1\}$, is the agents control process, which tells the market-maker whether or not to be posted at the best bid/ask.
-  • $Q^{\delta}=(Q_{t}^{\delta})_{\{0\leq t\leq T\}}$ is the agents’ controlled inventory process and is impacted by how much the agent trades.
-  • $S^{\delta}=(S_{t}^{\delta})_{\{0\leq t\leq T\}}$ is the controlled midprice process of the financial asset being traded, which is also impacted by how much the agent trades.
+What makes this particular trading problem special is that it begins to incorporate adverse selection in the trading process, which is a big hurdle for a MM in reality.
+Firstly, we would like to introduce the MM problem studied, as given in Cartea et al., (2015), where we also introduce an additional probability measure for non-adverse fills.
+We begin by describing the key stochastic processes as follows,
+  • $\delta^{\pm}=(\delta_{t}^{\pm})_{0\leq t\leq T}$, where $\delta_{t}^{\pm}\in\{0,1\}$, represents the agent’s control process.
+This variable determines whether the MM is posted at the best bid/ask.
+In our setting, each posted unit is of size 1.
+  • $\rho^{\delta}=(\rho^{\delta}_{t})_{0\leq t\leq T}$ denotes the probability that a posted limit order will be filled, where the agent has one unit posted at the best bid/ask.
+As in Cartea et al., (2015) and our reasoning given above, it is assumed that all trade order fills in this problem are non-adverse.
+However, the improved simulation framework later on will include adverse fills to show how performance can be significantly affected.
+Due to the dimension reduction that is needed and shown later on, adverse fills cannot be included in the problem setup.
+We highlight why this shortcoming occurs later.
+  • $Q^{\delta}=(Q_{t}^{\delta})_{\{0\leq t\leq T\}}$ is the agents’ controlled inventory process and is impacted by the amount the agent trades.
+  • $S^{\delta}=(S_{t}^{\delta})_{\{0\leq t\leq T\}}$ is the controlled midprice process of the financial asset being traded, which is also impacted by the amount the agent trades.
   • $M^{\pm}=(M_{t}^{\pm})_{\{0\leq t\leq T\}}$ is the MO arrival process.
@@
   • The agents cash process satisfies the SDE,
-    $dC_{t}^{\delta}=\displaystyle{\left(S_{t}+\frac{\Delta}{2}\right)}dN_{t}^{\delta,+}-\displaystyle{\left(S_{t}-\frac{\Delta}{2}\right)}dN_{t}^{\delta,-}$ (4)
+    $dC_{t}^{\delta}=\displaystyle{\left(S_{t}+\frac{\Delta}{2}\right)}dN_{t}^{\delta,+}-\displaystyle{\left(S_{t}-\frac{\Delta}{2}\right)}dN_{t}^{\delta,-}.$ (4)
     Here, $\Delta$ is the spread between the best bid and ask, and, as before, $N_{t}^{\delta,\pm}$ represents the counting process for filled LOs.
@@
 where $\mathscr{A}$ is the set of admissible strategies in which $\delta>0$ and uniformly bounded from above.
-It is also important to note that there is an inventory constraint in this problem, whereby the inventory process has an upper $\overline{q}$ and lower $\underline{q}$ bound.
-This essentially sets the maximum position the market maker can take, long or short.
+It is also important to note that there is an inventory constraint in this problem, by which the inventory process has an upper $\overline{q}$ and lower $\underline{q}$ bound.
+This essentially sets the maximum position that the MM can take long or short.
 Then, by applying the Dynamic Programming Principle (DPP), as explained in Cartea et al., (2015), the value function should satisfy the following Dynamic Programming Equation (DPE),
-$\begin{split}0=\left(\partial_{t}+\alpha\partial_{s}+\frac{1}{2}\sigma^{2}\partial_{ss}-\zeta\alpha\partial_{\alpha}+\frac{1}{2}\eta^{2}\partial_{\alpha\alpha}\right)H-\phi q^{2}\\ +\lambda^{+}\max_{\delta^{+}\in\{0,1\}}\{\mathbb{I}_{q>\underline{q}}\mathbb{E}\left[H(t,c+\left(S+\frac{\Delta}{2}\delta^{+}\right)\delta^{+},S,\alpha+\epsilon^{+},q-\delta^{+})-H\right]\\ +(1-\delta^{+}\mathbb{I}_{q>\underline{q}})\mathbb{E}\left[H(t,c,S,\alpha+\epsilon^{+},q)-H\right]\}\\ +\lambda^{-}\max_{\delta^{-}\in\{0,1\}}\{\mathbb{I}_{q<\overline{q}}\mathbb{E}\left[H(t,c-\left(S-\frac{\Delta}{2}\delta^{-}\right)\delta^{-},S,\alpha-\epsilon^{-},q+\delta^{-})-H\right]\\ +(1-\delta^{-}\mathbb{I}_{q<\overline{q}})\mathbb{E}[H(t,c,S,\alpha-\epsilon^{-},q)-H]\}\end{split}$ (7)
+$\begin{split}0&=\left(\partial_{t}+\alpha\partial_{s}+\frac{1}{2}\sigma^{2}\partial_{ss}-\zeta\alpha\partial_{\alpha}+\frac{1}{2}\eta^{2}\partial_{\alpha\alpha}\right)H-\phi q^{2}\\ &\quad+\lambda^{+}\bigg{(}\max_{\delta^{+}\in\{0,1\}}\bigg{\{}\rho(\delta^{+})\mathbb{I}_{q>\underline{q}}\mathbb{E}\Big{[}\delta^{+}\frac{\Delta}{2}+H(t,c+(S+\frac{\Delta}{2}\delta^{+})\delta^{+},S,\alpha+\epsilon^{+},\\ &\qquad q-\delta^{+})-H\Big{]}\bigg{\}}+\mathbb{E}\Big{[}H(t,c,S,\alpha+\epsilon^{+},q)-H\Big{]}\bigg{)}\\ &\quad+\lambda^{-}\bigg{(}\max_{\delta^{-}\in\{0,1\}}\bigg{\{}\rho(\delta^{-})\mathbb{I}_{q<\overline{q}}\mathbb{E}\Big{[}\delta^{-}\frac{\Delta}{2}+H(t,c-(S-\frac{\Delta}{2}\delta^{-})\delta^{-},S,\alpha-\epsilon^{-},\\ &\qquad q+\delta^{-})-H\Big{]}\bigg{\}}+\mathbb{E}\Big{[}H(t,c,S,\alpha-\epsilon^{-},q)-H\Big{]}\bigg{)},\end{split}$ (7)
 subject to the terminal and boundary conditions:
@@
   • Line 1: Characterizes the drift and diffusive parts of the midprice and short-term alpha processes, along alpha’s mean-reverting component.
-The last term on this line includes the running inventory penalty component, which only effects the inventory process.
-  • Line 2: Characterizes the change in the value function if the agent has a sell LO posted and the effect of an arriving buy MO filling the agent’s posted sell LO.
+The last term on this line includes the running inventory penalty component, which only affects the inventory process.
+  • Line 2: Characterizes the change in the value function if the agent has a sell LO posted and the effect of an arriving buy MO filling the agent’s posted sell LO, where it is clear that the midprice is unchanged, thus this represents a non-adverse trade order fill.
 This filled LO creates a jump in the short-term alpha component.
-The maximisation term characterizes the agent’s control, which determines whether to post a LO at the best offer.
+The maximization term characterizes the agent’s control, which determines whether to post an LO at the best offer.
   • Line 3: Characterizes the change in the value function when an MO arrives but the agent is not posted, in which case only the short-term alpha jumps.
-  • Line 4-5: Analogous to lines 2-3 but for the buy side of the LOB.
-In Cartea et al., (2015), they use an ansatz to simplify the DPE in equation (7).
-This ansatz splits out the accumulated cash, the book value of the shares which are marked-to-market at the midprice and the added value from making markets optimally as follows:
+  • Lines 4-5: Analogous to lines 2-3 but for the buy side of the LOB.
+In Cartea et al., (2015), they use an ansatz to simplify the DPE in Equation (7) to find a tractable solution.
+This ansatz splits out the accumulated cash, the book value of the shares which are marked-to-market at the midprice, and the added value from making markets optimally as follows:
 $H(t,c,S,\alpha,q)=c+qS+h(t,\alpha,q)$ (9)
 Substituting this into the DPE in Equation (7) gives us,
-$\begin{split}0=\displaystyle\left(\partial_{t}-\zeta\alpha\partial_{\alpha}+\frac{1}{2}\eta^{2}\partial_{\alpha\alpha}\right)h+\alpha q-\phi q^{2}\\ +\lambda^{+}\max_{\delta^{+}\in\{0,1\}}\displaystyle\left\{\mathbb{I}_{q>\underline{q}}\mathbb{E}\displaystyle\left[\delta^{+}\frac{\Delta}{2}+h(t,\alpha+\epsilon^{+},q-\delta^{+})-h(t,\alpha+\epsilon^{+},q)\right]\right\}\\ +\lambda^{-}\max_{\delta^{-}\in\{0,1\}}\displaystyle\left\{\mathbb{I}_{q<\overline{q}}\mathbb{E}\displaystyle\left[\delta^{-}\frac{\Delta}{2}+h(t,\alpha-\epsilon^{-},q+\delta^{-})-h(t,\alpha-\epsilon^{-},q)\right]\right\}\\ +\lambda^{+}\mathbb{E}[h(t,\alpha+\epsilon^{+},q)-h(t,\alpha,q)]\\ +\lambda^{-}\mathbb{E}[h(t,\alpha-\epsilon^{-},q)-h(t,\alpha,q)]\end{split}$ (10)
-subject to the terminal condition,
+$\begin{split}0&=\left(\partial_{t}-\zeta\alpha\partial_{\alpha}+\frac{1}{2}\eta^{2}\partial_{\alpha\alpha}\right)h+\alpha q-\phi q^{2}\\ &\quad+\lambda^{+}\bigg{(}\max_{\delta^{+}\in\{0,1\}}\bigg{\{}\rho(\delta^{+})\mathbb{I}_{q>\underline{q}}\mathbb{E}\Big{[}\delta^{+}\frac{\Delta}{2}+h(t,\alpha+\epsilon^{+},q)\Big{]}\\ &\qquad-h(t,\alpha+\epsilon^{+},q)\Big{]}\bigg{\}}+\mathbb{E}\Big{[}h(t,\alpha+\epsilon^{+},q)-h(t,\alpha,q)\Big{]}\bigg{)}\\ &\quad+\lambda^{-}\bigg{(}\max_{\delta^{-}\in\{0,1\}}\bigg{\{}\rho(\delta^{-})\mathbb{I}_{q<\overline{q}}\mathbb{E}\Big{[}\delta^{-}\frac{\Delta}{2}+h(t,\alpha-\epsilon^{-},q)\Big{]}\\ &\qquad-h(t,\alpha-\epsilon^{-},q)\Big{]}\bigg{\}}+\mathbb{E}\Big{[}h(t,\alpha-\epsilon^{-},q)-h(t,\alpha,q)\Big{]}\bigg{)}\end{split}$ (10)
+subject to the terminal condition:
 $h(T,\alpha,q)=-q\displaystyle\left(\frac{\Delta}{2}+\varphi q\right)$ (11)
-Lastly, by letting $\delta=0$, the optimal postings can be represented in a compact form due to the structure of the maximization terms as follows,
-$\begin{split}\delta^{+,*}(t,\alpha,q)=\mathbb{I}_{\{\frac{\Delta}{2}+\mathbb{E}[h(t,\alpha+\epsilon^{+},q-1)-h(t,\alpha+\epsilon^{+},q)]>0\}\cap\{q>\underline{q}\}}\\ \delta^{-,*}(t,\alpha,q)=\mathbb{I}_{\{\frac{\Delta}{2}+\mathbb{E}[h(t,\alpha-\epsilon^{-},q+1)-h(t,\alpha-\epsilon^{-},q)]>0\}\cap\{q<\overline{q}\}}\\ \end{split}$ (12)
-Here, we can see that the decision to post depends on the half-spread plus the expectation over the value functions computed at $\alpha\pm\epsilon^{\pm}$, where an arriving MO leads to a positive/negative jump in $\alpha$.
+As previously highlighted, a major shortcoming of this model is that adverse fills can no longer be directly incorporated into the solution.
+This occurs because the state variable $S$ has been completely eliminated from the problem through the ansatz in Equation (9).
+Since assessing adverse fills requires determining whether a price change has occurred, the absence of $S$ makes this evaluation impossible.
+However, in Section 5, we incorporate adverse fills into our improved simulation environment, allowing us to analyze the impact of their exclusion from the model.
+Lastly, by setting $\delta=0$ and assuming $\rho(\delta^{\pm})$ is either a constant or a function independent of the state variables, where for simplicity we assume that $\rho_{t}^{\delta}=\rho$ with $\rho\in(0,1)$, the optimal posting decisions can be expressed in a compact form due to the structure of the maximization terms:
+$\begin{split}\delta^{+,*}(t,\alpha,q)=\mathbb{I}_{\left\{\frac{\Delta}{2}+\rho\mathbb{E}[h(t,\alpha+\epsilon^{+},q-1)-h(t,\alpha+\epsilon^{+},q)]>0\right\}\cap\{q>\underline{q}\}}\\ \delta^{-,*}(t,\alpha,q)=\mathbb{I}_{\left\{\frac{\Delta}{2}+\rho\mathbb{E}[h(t,\alpha-\epsilon^{-},q+1)-h(t,\alpha-\epsilon^{-},q)]>0\right\}\cap\{q<\overline{q}\}}.\\ \end{split}$ (12)
+Here, we can see that the decision to post depends on the half-spread plus the expectation over the value functions computed at $\alpha\pm\epsilon^{\pm}$ weighted by a fill probability, where an arriving MO leads to a positive or negative jump in $\alpha$.
+This result is almost identical to the case in Cartea et al., (2015), except now with a non-adverse fill probability, $\rho$.
+This probability term can be interpreted as follows:
+  • If $10<\rho<$ 1, the threshold for placing an order is more restrictive, which means that the agent is less likely to post an LO because the expected benefit could be reduced.
+  • If $\rho=1$, this reduces to the original case in Cartea et al., (2015), which will be reflected in the benchmark environment. The optimal controls in Equation (12) remain the same.
 5 Trading Simulation Environment
-In this section we discuss the trading simulation environment used to test the performance of the optimal MM strategy.
+In this section, we discuss the trading simulation environment used to test the performance of the optimal MM strategy.
 In Cartea et al., (2015) and Jaimungal, (2019), they create a simulation environment in order to back-test this strategy, which we used as our benchmark.
-We will make one update to the benchmark by using real LOB data, rather than simulated prices.
+We will make one update to the benchmark using real LOB data, rather than simulated prices.
 This enabled us to back-test the strategy on real financial futures contracts data later on.
 We will discuss the extensions we made in our improved simulation environment, which we believe more accurately portrays how this type of MM strategy would perform in reality, albeit, still with some limitations.
-Specifically, we will describe how our framework tracks the adverse fills, as well as an improved fill probability structure related to non-adverse fills.
+Specifically, we will describe how our framework tracks adverse fills, as well as an improved fill probability structure related to non-adverse fills.
 We will then proceed to show how this can have a major effect on the performance of the trading strategy.
@@
 The products we will focus on include the near-term expiry contracts for ES, NQ, CL and ZN.
-Market makers often tend to focus on the more liquid products like these, in particular for HFT, as it will enable them to trade more volume.
+MMs often tend to focus on the more liquid products like these, in particular for HFT, as it will enable them to trade more volume.
 We resampled our price data into a 1 second time interval, which was then subsequently used as the time step size in our discrete environment.
+The data includes up to five levels of LOB events, with each new event generating a new data point at the next timestep.
+When resampling the dataset into 1-second intervals, we used the standard forward fill approach which ensures continuity by propagating the last observed value into missing time slots.
+Since market data arrives at irregular intervals, resampling often creates gaps where no new data exists.
+Forward fill addresses this by carrying forward the most recent data points until an update occurs, preserving the latest market state.
   • LO postings: Here the decision/control variable is stored, which tells us whether to be posted at the best bid/ask at any given time step.
@@
 See Jaimungal, (2019) for how this problem is solved numerically.
-  • Trade order fills: The benchmark trading simulation environment essentially assumes all fills are non-adverse, unless in the unlikely event a fill occurs at the exact same time as an adverse price movement.
-In our improved environment, however, we include adverse fills, as well as non-adverse fills, in the fill tracking process.
-The benchmark environment also assumes that all trade orders go to the front of the LOB queue, thus, their LOs will be the first to be filled.
-In most liquid financial assets, this is not very likely to occur.
-Based on this and using the empirical evidence from Tables 1 and 2 earlier, we will attach a binomial probability $p$ to all non-adverse fills.
-And so, non-adverse fills can be defined as follows,
-    $NFA_{t}=\sum_{i=1}^{N}\delta^{+}_{t_{i}}\mathbb{I}_{\{M^{+}_{t_{i}}=1\}}*p$ (13)
-    and
-    $NFB_{t}=\sum_{i=1}^{N}\delta^{-}_{t_{i}}\mathbb{I}_{\{M^{-}_{t_{i}}=1\}}*p,$ (14)
-    for $i=0,...,N$, where $N$ is the total number of time steps.
-Here, $NFA_{t}$ and $NFB_{t}$ are the counting processes for all non-adverse fills on the best ask and best bid, respectively.
-$p$ can be defined as:
-    $p=P(N_{t_{i}}^{\delta,\pm}|\delta_{t_{i}}^{+}\mathbb{I}_{\{AS(t_{i})\geq AS(t_{i+1})\}}=0,\delta_{t_{i}}^{-}\mathbb{I}_{\{BS(t_{i})\leq BS(t_{i+1})\}}=0)$ (15)
-    which is the probability of getting filled given the fill is non adverse.
-In Cartea et al., (2015) and Jaimungal, (2019), although not explicitly defined, this probability would be one, i.e., $p=1$, and thus we keep it like this in the benchmark environment.
-Next, we will keep track of the adverse fills in the following manner:
-    $AFA_{t}=\sum_{i=1}^{N}\delta_{t_{i}}^{+}\mathbb{I}_{\{AS(t_{i})<AS(t_{i+1})\}},$ (16)
-    and
-    $AFB_{t}=\sum_{i=1}^{N}\delta_{t_{i}}^{-}\mathbb{I}_{\{BS(t_{i})>BS(t_{i+1})\}},$ (17)
-    for $i=0,...,N$, where $AFA_{t}$ and $AFB_{t}$ are counting processes for adverse fills on the ask and adverse fills on the bid, respectively.
-Here, $BS_{t_{i}}$ and $AS_{t_{i}}$ represent the bid and ask prices, respectively.
+  • Trade order fills: The benchmark trading simulation environment assumes that all fills are non-adverse, except in the unlikely event that a fill occurs simultaneously with an adverse price movement.
+In our improved environment, however, we explicitly account for both adverse and non-adverse fills in the fill-tracking process.
+The benchmark environment also assumes that all trade orders are placed at the front of the LOB queue, meaning they are the first to be filled.
+In most liquid financial markets, this assumption is unrealistic.
+Based on this observation and empirical evidence from Tables 1 and 2, we introduced a probability function $\rho$ in Section 4 that governs the likelihood of non-adverse fills.
+And so, non-adverse fills can be defined as:
+    $NFA_{t}$ $=\sum_{t_{i}\leq t}\delta^{+}_{t_{i}}\mathbb{I}_{\{M^{+}_{t_{i}}=1\}}\rho$ (13)
+    $NFB_{t}$ $=\sum_{t_{i}\leq t}\delta^{-}_{t_{i}}\mathbb{I}_{\{M^{-}_{t_{i}}=1\}}\rho,$ (14)
+    where $NFA_{t}$ and $NFB_{t}$ are counting processes for all non-adverse fills at the best ask and best bid up to and including time $t$.
+Here, recall $\rho$ represents the probability of an order being filled at time $t_{i}$, given that it is posted at the best bid/ask and is a non-adverse fill.
+Unlike the benchmark environment where $\rho=1$, in reality, $\rho$ is influenced by a number of metrics such as LOB dynamics, queue position, and MO arrival rates, where we will use a rule of thumb to account for this based on the empirical evidence shown in Section 3.2.
+For adverse fills, we track them as follows:
+    $AFA_{t}$ $=\sum_{t_{i}\leq t}\delta_{t_{i}}^{+}\mathbb{I}_{\{AS(t_{i})<AS(t_{i+1})\}}$ (15)
+    $AFB_{t}$ $=\sum_{t_{i}\leq t}\delta_{t_{i}}^{-}\mathbb{I}_{\{BS(t_{i})>BS(t_{i+1})\}}$ (16)
+    where $AFA_{t}$ and $AFB_{t}$ are counting processes for adverse fills at the ask and bid up to and including time $t$.
+Here, $BS_{t_{i}}$ and $AS_{t_{i}}$ represent the bid and ask prices.
 To understand this intuitively, at each point in time, the trading strategy has a LO posted on the best bid (ask) $\delta_{t_{i}}^{-}(\delta_{t_{i}}^{+})$ and the asset price at time step $t_{i+1}$ is lower (higher) than at time $t_{i}$, the traders LO was then filled at the price $BS_{t_{i}}(AS_{t_{i}})$.
@@
 All other fills the trader receives will be considered non-adverse fills.
-Finally, in order to determine whether a fill occurred or not, we combine the two fill tracking formulas as follows,
-    $N_{t}^{\delta,+}=\sum_{i=1}^{N}max(AFA_{t_{i}},NFA_{t_{i}})$ (18)
-    and
-    $N_{t}^{\delta,-}=\sum_{i=1}^{N}max(AFB_{t_{i}},NFB_{t_{i}})$ (19)
-    where $N_{t}^{\delta,+}$ and $N_{t}^{\delta,-}$ are now counting processes collecting all the trade order fills that occur throughout this particular trading strategy.
+To ensure consistency in tracking all fills, we combine the non-adverse and adverse fill processes as follows:
+    $N_{t}^{\delta,+}$ $=\sum_{t_{i}\leq t}\left(AFA_{t_{i}}+NFA_{t_{i}}\right)$ (17)
+    $N_{t}^{\delta,-}$ $=\sum_{t_{i}\leq t}\left(AFB_{t_{i}}+NFB_{t_{i}}\right)$ (18)
+    where $N_{t}^{\delta,+}$ and $N_{t}^{\delta,-}$ are counting processes collecting all trade order fills occurring up to and including time $t$ under this trading strategy.
 For the processes involving MO arrivals, inventory, short-term alpha and cash, we follow the method given in Jaimungal, (2019), where they essentially discretize the Equations (2)-(4), as well as devising a unique way to simulate MOs.
-5 Results
-Here, we will discuss the results from testing the SOC MM strategy in our improved simulation environment and we will compare this to the benchmark environment.
-First, we would like to briefly summarize the numerical solution under our parameter set and we provide a visual in Figure 2, which looks similar to the solution given in Cartea et al., (2015) and Jaimungal, (2019).
+6 Simulation Results
+Here, we will discuss the results from testing the SOC MM strategy in our improved simulation environment and compare this with the benchmark environment.
+First, we would like to briefly summarize the numerical solution under our parameter set and provide a visual in Figure 3, which looks similar to the solution given in Cartea et al., (2015) and Jaimungal, (2019).
+We compare the solutions of the buy and sell side posts, where the top plots correspond to $\rho=1$ and the bottom plots to $\rho=0.2$, and these values were chosen as these specific solutions will be utilized in our strategy simulations in Section 5.
 One can see how the optimal solution varies as a function of time ($t$), short-term alpha ($\alpha$) and inventory($q$).
-The left (right) Figure portrays a visual of LOs posted on the best offer (bid), where a sell (buy) LO is posted if the MMs inventory is above (below) the surface.
-The images look almost identical as the arrival rates of MOs are the same on both sides of the LOB.
-Far from maturity, the strategy is mostly independent of time, but one can see that as the MM approaches the maturity time ($T$), the optimal strategy is essentially independent of the short-term alpha ($\alpha$), since the MM is solely focused on liquidating any remaining position.
-Figure 2: Optimal solution as a function of time, short-term alpha and inventory.
-As discussed earlier, the main difference in our improved simulation environment is that adverse fills are now included in the strategy simulation process, determined based on equations (16)-(17).
-Our aim in this section is to show how adverse fills and fill probabilities on non-adverse fills can significantly effect the performance of the trading strategy.
+The left (right) figure displays a visual of LOs posted on the best offer (bid), where a sell (buy) LO is posted if the MMs inventory is above (below) the surface.
+It is clear that a lower non-adverse fill probability, $\rho$, leads to a slightly different solution where the agent is less likely to post LOs at lower inventory levels.
+The top left and top right plots, as well as the bottom left and bottom right plots, appear almost identical due to the equal arrival rates of MOs on both sides of the LOB.
+Far from maturity, the strategy is mostly independent of time, but one can see that as the MM approaches the maturity time ($T$), the optimal strategy is essentially independent of the short-term alpha ($\alpha$), since the MM is focused solely on liquidating any remaining position.
+As discussed earlier, the main difference in our improved simulation environment is that adverse fills are now included in the strategy simulation process, determined according to Equation (16).
+Our aim in this section is to show how adverse fills and fill probabilities on non-adverse fills can significantly affect the performance of the trading strategy.
 We will discuss the results for CL in depth here, where the same repeated analysis for ES, NQ and ZN can be found in the appendix.
-Based on the empirical evidence for the distribution between adverse and non-adverse fills given in Table 2, we also set the non-adverse fill probability to $p=0.2$.
+Based on the empirical evidence for the distribution between adverse and non-adverse fills given in Table 2, we also set the non-adverse fill probability at $\rho=0.2$.
 See Table 4, below, for a list of the parameter values used.
-We kept most of the parameters similar to Cartea et al., (2015) and Jaimungal, (2019), however, we decreased the time step and increased the maturity time, in order to make the length of time more realistic to the data.
-We also decreased the intensity of the MOs as another way to bring the distribution of adverse and non-adverse fills a bit closer to reality.
-A few slight adjustments to some of the other parameters were also needed to keep the numerical solution stable.
+We kept most of the parameters similar to Cartea et al., (2015) and Jaimungal, (2019), however, we decreased the time step and increased the maturity time, to make the length of time more realistic to the data.
+Some slight adjustments to some of the other parameters were also needed to keep the numerical solution stable.
+Figure 3: Optimal solution as a function of time, short-term alpha, and inventory.
+The top plots correspond to $\rho=1$, while the bottom plots correspond to $\rho=0.2$.
+Here, darker red indicates a larger long position, while darker blue indicates a larger short position.
+Now, we will review the strategy simulation from our improved simulation environment alongside the performance of the benchmark environment.
+In Figure 4, we begin by showing a snapshot of the strategy over a random 120 second path in CL, where one can see when the MM is posted on the best bid/ask and when they receive trade order fills.
+The left and right figures show the strategy simulation in the benchmark and improved environments, respectively.
+Here, the green (blue) lines indicate when the MM is posted on the best bid (ask), the filled circles indicate when the MMs LO is filled on the best bid/ask, and the unfilled circles indicate times the MM would have been filled if they had a LO posted.
+One noticeable feature, in the figure on the left, is that whenever the agent is posted and the price moves through their order, they do not automatically receive a fill.
+This, as we mentioned earlier, is because MOs are simulated independently from the price process in the benchmark models in Cartea et al., (2015) and Jaimungal, (2019), and is contrary to what would happen in reality.
+And so, all the fills in this left figure would be referred to as non-adverse fills.
+In the right figure, one can see where all the adverse fills would have occurred (denoted by AFB and AFA).
+Note that this can also alter the posting strategy later on because the inventory process now evolves differently to how it would in the benchmark environment.
 Parameters
@@
 $\varphi$ 0.01 $\phi$ 0
-$\lambda^{+}$ 0.1458 $\lambda^{-}$ 0.1458
+$\lambda^{+}$ 0.5833 $\lambda^{-}$ 0.5833
 Ndt 120 dt 1 second
 Table 3: Simulation parameters.
-Now, we will review the strategy simulation from our improved simulation environment alongside the performance of the benchmark environment.
-In Figure 3, we begin by showing a snapshot of the strategy over a random 120 second path in CL, where one can see when the market-maker is posted on the best bid/ask and when they receive trade order fills.
-The left and right Figures show the strategy simulation in the benchmark and improved environments, respectively.
-Here, the green (blue) lines indicate when the market-maker is posted on the best bid (ask), the filled circles indicate when the market-makers LO is filled on the best bid/ask, and the unfilled circles indicate times the market-maker would have been filled if they had a LO posted.
-One noticeable feature, in the left Figure, is that whenever the agent is posted and price moves through their order, they do not automatically receive a fill.
-This, as we mentioned earlier, is because MOs are simulated independently from the price process in Cartea et al., (2015), Cartea et al., 2018b and Jaimungal, (2019), and is contrary to what would happen in reality.
-And so, all the fills in this left Figure would be referred to as non-adverse fills.
-In the right Figure, one can see where all the adverse fills would have occurred (denoted by AFB and AFA).
-Note, this can also alter the posting strategy later on because the inventory process now evolves differently to how it would in the benchmark environment.
-Figure 3: A snapshot of a random path of the strategy in the benchmark (left) and improved (right) simulation environments.
-The top line (black/blue) indicates the best ask, the middle line (black) the midprice and the bottom line (black/green) the best bid.
-When the top (bottom) line is blue (green), the agent is posted on the best ask (bid).
-Closed circles indicate fills received when the market-maker is posted and open circles indicate fills that the market-maker would’ve received if they were posted.
-It’s quite apparent that the timing of the trade order fills has changed and many trade order fills occur at very unfavourable prices.
-Earlier, in our simple trading example, where we gave results in Table 2, we saw there that a overwhelming majority of the fills for the strategy simulation in CL were adverse, as well as for ES, NQ and ZN.
-This test was also conducted on the exact same trading day as this simulation.
-In Table 4, we show the number of each type of fill that occurred within our improved simulation environment.
-We can see that using a much lower non-adverse fill probability of $p=0.2$ created a similar distribution of fills between non-adverse fills and adverse fills as in Table 2, and we will soon see that the overall performance has now dropped significantly.
-This trend also persists in the strategy simulations using ES, NQ and ZN data, as can be seen in the appendix.
-Based on our empirical evidence, we believe these results more accurately portray how the strategy would perform in reality.
-However, we may not want to be overly pessimistic by thinking the SOC MM strategy should have the same number of adverse fills as the basic posting strategy, because the SOC MM model has a short-term alpha predictor which could potentially predict some of these adverse fills.
-However, it’s extremely unlikely that any short-term alpha predictor could foresee all future adverse fills.
+Figure 4: A snapshot of a random strategy path in the benchmark (left) and improved (right) simulation environments.
+The top (black/blue), middle (black), and bottom (black/green) lines represent the best ask, midprice, and best bid, respectively.
+Blue (green) segments indicate the agent is posted on the best ask (bid).
+Closed circles show fills received, while open circles represent fills the MM would have received if posted.
 Fill Type Amount
@@
 NFB 813
-Table 4: Number of trade order fills that were adverse and non-adverse.
-Next, in Figure 4, one can see the wealth and inventory path for the same random 120 second run of the strategy.
-Here, again, the left and right figures are under the benchmark and improved simulation environments, respectively.
+Table 4: Number of trade order fills that were adverse and non-adverse over 330 simulations throughout the trading on April 24th, 2025.
+It is quite apparent that the timing of the trade order fills has changed and many trade order fills occur at very unfavorable prices.
+Earlier, in our simple trading example, where we gave results in Table 2, we saw there that a overwhelming majority of the fills for the strategy simulation in CL were adverse, as well as for ES, NQ and ZN.
+This test was also conducted on the exact same trading day as this simulation.
+In Table 4, we show the number of each type of fill that occurred within our improved simulation environment.
+We can see that using a much lower non-adverse fill probability of $\rho=0.2$ created a similar distribution of fills between non-adverse fills and adverse fills as in Table 2, and we will soon see that the overall performance has now dropped significantly.
+This trend also persists in the strategy simulations using ES, NQ and ZN data, as can be seen in the appendix.
+Based on our empirical evidence, we believe that these results better depict how the strategy would perform in reality.
+However, we may not want to be overly pessimistic by thinking the SOC MM strategy should have the same number of adverse fills as the basic posting strategy, because the SOC MM model has a short-term alpha predictor, which could potentially predict some of these adverse fills.
+However, it is highly unlikely that any short-term alpha predictor could anticipate all future adverse fills.
+The simple predictor in Equation (2) lacks the sophistication needed for such predictions, especially given its simplicity and the likely alpha decay it has undergone due to its public availability.
+Recall our earlier note on the alpha decay of a publicly disclosed trading strategy in Section 3.2.
+Next, in Figure 5, one can see the wealth and inventory path for the same random 120-second run of the strategy.
+Here, again, the left and right figures are in the benchmark and improved simulation environments, respectively.
 The wealth path is marked-to-market at each time step as in Equation (4) and the inventory path changes whenever a trade order fill is received as in Equation (3).
-Here, the market-makers wealth path in the left Figure finishes positive, and in the right Figure it finishes negative, over the same 120 second path.
-The inventory paths’ show that the benchmark environment traded a lot less than our improved environment, which is mostly due to the additional adverse fills this trader received.
-The additional adverse fills also significantly outweighed the reduced number of non-adverse fills.
-And so, it’s quite apparent that these paths now evolve very differently.
-Lower non-adverse fill probabilities and adverse fills has significantly effected how these processes evolved.
+Here, the MM’s wealth path in the left figure ends positive and at a higher value than in the right figure, both over the same 120 second price path.
+The inventory paths show that the benchmark environment traded quite differently from our improved environment, which is largely due to the additional adverse fills this trader received, but also the lower non-adverse fill probability.
+Thus, it is quite apparent that these paths now evolve very differently.
+Lower non-adverse fill probabilities and adverse fills have significantly affected how these processes evolved.
+Figure 5: Wealth and inventory processes for the same random path of the strategy under the benchmark (left) and improved (right) simulation environments.
+Figure 6: Terminal cash histogram for the benchmark (left) and improved (right) simulation environments over all 330 simulation paths.
 In order to tell the full story, we show a histogram, in Figure 6, of the terminal cash values from all 330 simulation runs of the strategy over this trading day.
@@
 One can see that, overall, the strategy performed reasonably well in the benchmark environment and much worse in our improved environment.
-We believe the performance metric in the benchmark environment is over-inflated, as the market-maker received a lot more non-adverse fills than likely in reality, as well not receiving any adverse fills, which clearly significantly effects the performance.
-In the right Figure, we can see that the strategy does not perform as well and regularly suffers large losses, relatively speaking.
-This confirms to us that simulating the performance of a short-term style trading strategy, in particular one involving posting many LOs, is very misleading if it does not track adverse fills and include more accurate non-adverse fill probabilities.
+We believe the performance metric in the benchmark environment is over-inflated, as the MM received a lot more non-adverse fills than likely in reality, as well not receiving any adverse fills, which clearly significantly affects the performance.
+This confirms to us that simulating the performance of a short-term style trading strategy, in particular one that involves posting many LOs, is very misleading if it does not track adverse fills and include more accurate non-adverse fill probabilities.
 This is just as evident in ES, NQ and ZN, where similar results for these assets can be seen in the appendix.
-Figure 4: Wealth and inventory processes for the same random path of the strategy under the benchmark (left) and improved (right) simulation environments, respectively.
-Figure 5: Terminal cash histogram for the benchmark (left) and improved (right) simulation environments over all 330 simulation paths.
 7 Conclusion and Future Recommendations
-In this paper we simulated the performance of a SOC optimal MM strategy that both included and excluded adverse fills, as well as adjusting the probability of a non-adverse fill.
-Many MM style strategies often trade passively by posting many LOs throughout the LOB, thus the results here are particularly relevant to these types of strategies, but also to any trading strategy that posts a large amount LOs.
-Models in mathematical finance often make simplifying assumptions in order to find robust solutions to a problem and while many don’t often effect the overall performance too much, some certainly can.
-In our setting, modelling the price process and the arrival of MOs independently can create large inaccuracies, which is clear from our simulation results.
-Specifically, our finding shows that excluding adverse fills and non-adverse fill probabilities can significantly inflate the performance of a trading strategy.
+In this paper we simulated the performance of an SOC optimal MM strategy that included and excluded adverse fills, as well as adjusting the probability of a non-adverse fill.
+Many MM style strategies often trade passively by posting many LOs throughout the LOB, thus the results here are particularly relevant to these types of strategies, but also to any trading strategy that posts a large amount of LOs.
+Models in mathematical finance often make simplifying assumptions in order to find robust solutions to a problem, and while many do not often affect the overall performance too much, some certainly can.
+In our setting, modeling the price process and the arrival of MOs independently can create large inaccuracies, which is clear from our simulation results.
+Specifically, our findings show that excluding adverse fills and non-adverse fill probabilities can significantly inflate the performance of a trading strategy.
 Adverse trade order fills are never completely avoidable for a trader who devises a LO posting strategy and are particularly apparent in short-term strategies.
-Although we can try to predict short-term price movements using methods like the short-term alpha predictor given in Cartea et al., (2015) and Cartea et al., 2018b , large orders often randomly enter the market taking out price levels on a whim.
+Although we can try to predict short-term price movements using methods like the short-term alpha predictor given in Cartea et al., (2015) and Cartea et al., 2018b , these methods are quite simplistic and large orders often randomly enter the market, taking out price levels on a whim.
 Thus, if a MM or trader is posted at these price levels they will receive adverse fills as there usually isn’t enough time to cancel the orders, regardless of some of the speed advantageous some HFTs may have.
-In regard to future research recommendations, we think the results could be improved if a more dynamic MM model were used, where the MM can update the important parameters over time.
-It is clear from our analysis and the literature that LOB data has a clustering effect, thus this could enable the MM to update their non-adverse fill probabilities, as well as improving the short-term alpha predictors ability to predict adverse moves.
-Papers like Arroyo et al., (2024) and Maglaras et al., (2022) are some new initial works that aim to estimate fill probabilities using deep learning approaches, which seem to develop some promising ideas for devising a more dynamic model.
-One should also assess the performance of any short-term alpha predictor used to predict adverse moves, as this is one of the key components of the model in this paper.
-In the SOC framework, this simple short-term alpha predictor helped bring tractable results, but it is unlikely to be that simple to make such price predictions in reality.
-Lastly, one could also try to improve the price process modelling aspect, by using a jump-diffusion or Levy-style model.
-These price processes have been proven to more accurately mimic LOB dynamics rather than the general diffusion models, although sometimes more difficult to apply directly in a SOC model.
+In terms of future research recommendations, we believe that the results could be improved if a more dynamic MM model was used, where the MM could update the important parameters over time.
+For instance, the model could simulate MOs entering the market with varying sizes, whereas we currently assume that each order has a size of 1.
+It is clear from our analysis and the literature that LOB data has a clustering effect, thus this could more accurately simulate adverse fills, as well as enabling the MM to update their non-adverse fill probabilities and perhaps improve the short-term alpha predictors ability to predict adverse moves.
+Another extension could be to adjust the posting strategy in the SOC problem, as discussed in Section 3.2, by spreading out LO postings more for assets that experience rapid price changes, such as NQ.
+Papers like Arroyo et al., (2024) and Maglaras et al., (2022) are some new initial studies that aim to estimate fill probabilities using deep learning approaches, which seem to develop some promising ideas for devising a more dynamic model.
+The performance of any short-term alpha predictor used to predict adverse moves should also be assessed, as this is one of the key components of the model in this paper.
+In the SOC framework, this simple short-term alpha predictor helped bring tractable results, but it is unlikely that it would be that simple to make such price predictions in reality.
+Lastly, one could also try to improve the price process modeling aspect, by using a jump-diffusion or Levy-style model.
+These price processes have been proven to more accurately mimic LOB dynamics rather than the general diffusion models, although they are sometimes more difficult to apply directly in a SOC model.
 Acknowledgements
@@
 Fill Type Amount Fill Type Amount Fill Type Amount
-AFA 5904 AFA 12319 AFA 478
-NFA 789 NFA 840 NFA 825
-AFB 5873 AFB 12370 AFB 468
-NFB 774 NFB 781 NFB 821
+AFA 5804 AFA 11509 AFA 389
+NFA 2107 NFA 1963 NFA 1960
+AFB 5780 AFB 11448 AFB 432
+NFB 2088 NFB 2009 NFB 1999
 Table 5: Number of fills received over each 120 second run of the strategy that were adverse and non-adverse in ES, NQ and ZN, from left to right.
+Note that the non-adverse fill probability should be much lower in ZN as queue sizes are much larger relative to CL, ES, and NQ while trade sizes are very similar, hence the different distribution and more inflated performance results.